### v2

# suppSupplementary References

Abstract
In multi-armed bandit (MAB) experiments, it is often advantageous to continuously produce inference on the average treatment effect (ATE) between arms as new data arrive and determine a data-driven stopping time for the experiment. We develop the Mixture Adaptive Design (MAD), a new experimental design for multi-armed bandit experiments that produces powerful and anytime-valid inference on the ATE for any bandit algorithm of the experimenter’s choice, even those without probabilistic treatment assignment. Intuitively, the MAD “mixes” any bandit algorithm of the experimenter’s choice with a Bernoulli design through a tuning parameter $\delta_{t}$, where $\delta_{t}$ is a deterministic sequence that decreases the priority placed on the Bernoulli design as the sample size grows. We prove that for $\delta_{t}=\omega\left(1/t^{1/4}\right)$, the MAD generates anytime-valid asymptotic confidence sequences that are guaranteed to shrink around the true ATE. Hence, the experimenter is guaranteed to detect a true non-zero treatment effect in finite time. Additionally, we prove that the regret of the MAD approaches that of its underlying bandit algorithm over time, and hence, incurs a relatively small loss in regret in return for powerful inferential guarantees. Finally, we conduct an extensive simulation study exhibiting that the MAD achieves finite-sample anytime validity and high power without significant losses in finite-sample reward.
Keywords: Adaptive Experimental Design, Multi-armed Bandit, Online Learning, Sequential Analysis, Always-valid inference, Asymptotic Confidence Sequence, A/B Test
1 Introduction
In many large-scale online experimentation settings, companies increasingly favor multi-armed bandit (MAB) experiments over traditional randomized control trials (i.e., A/B tests) to reduce the risk of experimentation (Kim and Lim 2016; Kohavi et al. 2020a; Zhang et al. 2020; Bojinov and Gupta 2022a; Aramayo et al. 2023). In MAB experiments, the learning algorithm assigns one of a set of treatments at each time point, where the treatment assignment probabilities can depend on previous data in the experiment, i.e., the outcomes and treatment assignments from prior time points. Since MABs adaptively update treatment assignment probabilities based on past data, they generally decrease the proportion of units exposed to sub-par treatments, hence mitigating potential adverse effects due to experimentation. Though MABs are primarily designed for risk minimization, managers and decision-makers often want to conduct statistical inference on the average treatment effect (ATE) between arms at the end of the MAB experiment, not just minimize risk. For instance, various large-scale analyses of online experiments have shown that only a relatively small percentage of product changes produce positive effects (${}20\%$) Kohavi et al. (2020b); Bojinov and Gupta (2022b). To pinpoint the limited set of modifications that deliver customer value while simultaneously protecting against changes that would lead to negative effects, managers must be able to conduct inference on experiments to measure these effects and assess their statistical significance. Such analyses of experiments enables organizational learning that can improve product features in the future.
By nature of their risk minimization, MAB algorithms generally undersample from subpar arms, making statistical inference challenging as observations from these subpar arms are necessary to evaluate and gain precise uncertainty quantification on the ATE. Various approaches have designed inferential methods to analyze MAB experiments at the end of the experiment Luedtke and Van Der Laan (2016); Deshpande et al. (2019); Hadad et al. (2021); Zhang et al. (2020, 2021); Nair and Janson (2023), and there is growing literature focused on developing experimental designs to increase power and ensure valid inference at the end of a MAB experiment with a pre-specified time horizon Kasy and Sautmann (2021); Simchi-Levi and Wang (2023), though they often focus on specific bandit algorithms to facilitate their theoretical analysis. However, since the fundamental idea behind MABs is to use existing data to inform future decisions, a more natural inferential framework would be to allow the experimenter to continuously produce inference on the average treatment effect (ATE) between arms as new data arrive, i.e., to perform anytime valid inference Waudby-Smith and Ramdas (2020); Howard et al. (2021); Johari et al. (2022); Waudby-Smith et al. (2022). Anytime valid inference enables the experimenter to stop an experiment at a data-driven stopping time e.g., when statistically significant results are achieved or if an outsized harmful effect is observed. Thus, experimenters can continuously perform statistical inference on their experiment, enabling proactive risk mitigation. However, existing works for anytime valid inference in adaptive experiments often places restrictive assumptions on the kind of adaptivity allowed Howard et al. (2021); Ham et al. (2023), hence excluding nearly all common bandit algorithms such as Thompson sampling Thompson (1933) and the Upper Confidence Bound algorithm Auer et al. (2002a).
In this work, we develop the Mixture Adaptive Design (MAD), a new experimental design for MAB experiments that generates powerful and anytime-valid inference on the ATE for any bandit algorithm. Intuitively, the MAD “mixes” a bandit algorithm with a Bernoulli design through a tuning parameter $\delta_{t}$, where $\delta_{t}\in(0,1]$ is a deterministic, i.e., non-random, sequence that controls the priority placed on the Bernoulli design as the sample size grows. For $\delta_{t}=\omega\left(\frac{1}{t^{1/4}}\right)$, we provide a confidence sequence that is asymptotically anytime valid and guaranteed to shrink around the true ATE for any choice of bandit algorithm. This confidence sequence is, to our knowledge, the first confidence sequence for ATE estimation in adaptively collected data settings that does not require the treatment assignment probabilities to be bounded away from zero and one. Thus, the MAD expands the utility of confidence sequences to nearly any bandit algorithm while providing guarantees on the validity and power of the ATE estimation. From a practical perspective, the Mixture Adaptive Design guarantees that, under a true non-zero treatment effect, an experimenter with the goal of stopping her experiment once the confidence sequence suggests a non-zero treatment effect is guaranteed to see statistically significant results in finite time. The condition that $\delta_{t}=\omega(\frac{1}{t^{1/4}})$ provides the user great flexibility to tune the experimental design based on the problem specifics. For instance, setting $\delta_{t}=1$ would recover a Bernoulli design. We provide recommendations for setting $\delta_{t}$ based on the user’s priorities; see Section 6. Finally, we show empirically that the MAD generates confidence sequences that shrink quickly around the true ATE and achieves the correct coverage in finite samples without major losses in reward compared to a standard bandit design.
2 Background and Related Work
Simultaneously conducting inference on the ATE while minimizing regret is particularly challenging because MABs determine treatment assignments adaptively; most statistical guarantees of ATE estimation methods for independent and identically distributed (i.i.d.) data no longer hold in the MAB setting. For instance, the difference in sample averages between a treatment arm and a control arm is no longer unbiased for the true ATE of these arms (Xu et al. 2013) and may no longer be asymptotically normal Dimakopoulou et al. (2021); Zhang et al. (2020). While Inverse Propensity Weighting (IPW) estimators are often unbiased in the MAB setting Horvitz and Thompson (1952); Hadad et al. (2021), their variance can explode, thus making any subsequent hypothesis testing on the ATE essentially powerless. The variance can rapidly increase as the probability that a bandit algorithm pulls a sub-optimal arm rapidly approaches zero (Russo et al. 2020). For instance, the well-known Thompson Sampling (TS) algorithm Agrawal and Goyal (2012) and variations of the Upper Confidence Bound (UCB) algorithm Garivier and Cappé (2011); Ménard and Garivier (2017); Kaufmann (2018) achieve a regret bound of $O(\log T)$ and are said to be asymptotically optimal as the regret of any algorithm must have at least a $\log(T)$ rate Lai and Robbins (1985). Hence, we expect the number of draws from sub-optimal arms to be on the order of $\log(T)/T$ in the long run Kasy and Sautmann (2021).
Prior works on experimental design for MAB experiments have primarily adapted existing bandit algorithms to sample more from sub-optimal arms, making them more amenable to ATE estimation. For instance, Kasy and Sautmann (2021) proposes a variation of Thompson sampling which re-weights the estimated mean reward of each arm to enforce additional exploration of sub-optimal arms. Simchi-Levi and Wang (2023) first formulates the trade-off between mean squared estimation error and regret via a minimax multi-objective optimization problem, then proposes a variation of the EXP3 algorithm Auer et al. (2002b) which takes draws from the estimated sub-optimal arms via a specified rate and shows that this adaptation satisfies this characterization of optimality. Intuitively, some degree of exploration must be imposed to produce inferential guarantees, as the precision of ATE inference essentially depends on the variance of the treatment arm which has the fewest number of observations. However, these works focus on adapting specific bandit algorithms and thus, are difficult to generalize and apply broadly to other bandit algorithms. Additionally, these works propose designs that do not allow the user to tune the degree of exploration in the algorithm Kasy and Sautmann (2021) or have limitations on when during the experiment this tuning can take place Simchi-Levi and Wang (2023). Hahn et al. (2011) proposes a two-stage adaptive design and proves an asymptotic normality result for ATE estimation in this setting, but requires the adaptive assignment algorithm to have assignment probabilities bounded away from zero and one, which is not satisfied by most common bandit algorithms such as UCB (a deterministic algorithm) and Thompson sampling. Other works such as Hadad et al. (2021); Dimakopoulou et al. (2018) utilizes doubly-robust estimators from causal inference literature to estimate the mean reward of each arm, but focus specifically on improving best arm identification within the bandit algorithm to produce better finite-sample regret rather than generating inference on the ATE. Works such as Lai and Wei (1982); Hadad et al. (2021); Luedtke and Van Der Laan (2016); Deshpande et al. (2019); Zhang et al. (2020, 2021); Nair and Janson (2023) develop statistical tests or CLTs for adaptively collected data settings without significantly altering the design of the bandit algorithm, but often requires certain assumptions on the outcome distributions, such as moment bounds or independently and identically distributed outcomes. Most importantly, all of the above approaches are designed for the setting where the experimenter is conducting inference at the end of an adaptive experiment with a pre-specified time horizon.
However, since the fundamental idea behind MABs and adaptive experiments in general is to use existing data to inform future decisions, a more natural inferential framework would be to allow the experimenter to continuously produce inference on the ATE between arms as new data arrive. Such a framework would allow experimenters to “peek” at the data in their experiment without invalidating the subsequent statistical inference. Therefore, experiments would not require the experimenter to specify a sample size in advance, but determine a data-dependent stopping time in response to running inferential results (e.g., stop the experiment once statistically significant results are achieved), further minimizing harm to experimental units and reducing the opportunity cost of running long experiments.
Such a continuous monitoring setting requires statistical tests that uniformly control Type I error at every time point, i.e., are anytime-valid. As classical tests do not satisfy this condition, prior works have proposed using confidence sequences to enable valid inference Darling and Robbins (1967); Lai (1976a, b); Johari et al. (2022); Waudby-Smith et al. (2023).
Definition 1 (Confidence Sequence).
A confidence sequence (CS) is a set of confidence sets $\{C_{t}\}_{t=1}^{\infty}$ at level $\alpha\in(0,1)$ such that, for a true (non-zero) treatment effect $\tau_{t}$,
$\mathbb{P}\left(\forall t,\tau_{t}\in C_{t}\right)\geq 1-\alpha.$
Johari et al. (2022) provides a non-asymptotic confidence sequence for traditional A/B testing settings where the treatment allocations are not made adaptively. However, extending non-asymptotic confidence sequences to adaptively collected data often require specific assumptions on the data such as known bounds on the random variables and/or tail behavior assumptions such as sub-Gaussianity Waudby-Smith and Ramdas (2020); Howard et al. (2021); Howard and Ramdas (2022). In particular, such assumptions prevent non-asymptotic CSs from being applicable or even possible to construct in many real-world settings. To address this restriction, Waudby-Smith et al. (2023) first introduced the notion of an asymptotic confidence sequence and derived a universal asymptotic CS that requires only CLT-like assumptions.
Definition 2 (Asymptotic Confidence Sequence).
$(\hat{\mu}_{t}\pm\hat{V}_{t})$ is an asymptotic $1-\alpha$ confidence sequence for a target parameter $\mu_{t}$ if there exists some (unknown) non-asymptotic $1-\alpha$ confidence sequence $(\hat{\mu}_{t}\pm V^{*}_{t})$ for $\mu_{t}$ such that
$\frac{\hat{V}_{t}}{V^{*}_{t}}\stackrel{{\scriptstyle a.s.}}{{\to}}1.$
Furthermore, we say that $\hat{V}_{t}$ has approximation rate $R_{t}$ if $V^{*}_{t}=\hat{V}_{t}+o(R_{t})$ almost surely, where $R_{t}$ is a deterministic (i.e., non-random) sequence.
The assumptions necessary for generating asymptotic CSs are comparatively much weaker, and hence, asymptotic CSs expand the utility of CSs to a wider array of settings. For instance, Ham et al. (2023) proposes a framework for continuous design-based causal inference for time series and panel experiments utilizing asymptotic CSs. However, existing work on continuous inference for MAB experiments is rather limited. While Ham et al. (2023) and Howard et al. (2021) discuss extensions to experiments with adaptive treatment assignment probabilities, their proposed CSs require probabilistic treatment assignment, i.e., they assume that there exists $20<p_{min}\leq 1/$ 2 s.t. $p_{t}(w)\in[p_{min},1-p_{min}]$ almost surely. This restriction excludes many commonly used bandit algorithms such as UCB, which has deterministic treatment assignments, and Thompson Sampling. Without this treatment assignment assumption, the asymptotic confidence sequences of Ham et al. (2023) are not guaranteed to be asymptotically anytime-valid Waudby-Smith et al. (2023), and the non-asymptotic confidence sequences of Howard et al. (2021) may be inapplicable since their validity results requires $p_{min}$ to exist and be known. Thus, applying such approaches to most commonly used MAB algorithms necessarily requires imposing lower bounds on the treatment assignment probabilities of the algorithm, such as clipping them to be within $[p_{min},1-p_{min}]$, restricting the adaptivity allowed and increasing regret.
3 Problem Statement
In this section, we formalize the multi-armed bandit problem setting and define the causal estimands of interest. Throughout, we adopt a design-based causal inference framework Neyman and Iwaszkiewicz (1935); Fisher (1936); Ding et al. (2016); Abadie et al. (2020); thus, instead of treating the outcomes as random variables, we condition on the set of potential outcomes and treat only the assignment as random. Our estimand of interest is the sample average treatment effect, rather than the population average treatment effect for some hypothetical super-population. This design-based framework allows us to make relatively few assumptions on the outcome distribution of our bandit algorithm, i.e., we do not assume any parametric form on the outcome distribution or that our outcomes are independent and identically distributed (i.i.d.) from some distribution Dimakopoulou et al. (2018); Zhang et al. (2020); Hadad et al. (2021); Banerjee et al. (2023). In many real-world settings, we expect that the outcomes will be dependent or non-stationary Besbes et al. (2014); Allesiardo et al. (2017); Wu et al. (2018). In such settings, incorrectly assuming i.i.d. outcomes can result in effective sample sizes that are much smaller than assumed and cause variance estimation to be highly conservative Meng (2018).
Assume that we observe a sequence of $t$ units $\{W_{i},Y_{i}\}_{i=1}^{t}$ where $W_{i}\in\{0,...,K-1\}$ and $Y_{i}$ are the treatment assignment and outcome respectively for unit $i$. In the MAB literature, $\{W_{i},Y_{i}\}_{i=1}^{t}$ are analogous to the sequence of actions and rewards commonly notated $\{A_{i},R_{i}\}_{i=1}^{t}$. Although we will ultimately show that our method can be defined more broadly to any $K\geq 2$ number of treatments and the batched bandit setting, we first assume that we observe a single unit at each time and that we have binary treatment assignments, i.e., $W_{i}\in\{0,1\}$.
In our binary treatment setting, each unit $i$ has a pair of potential outcomes $\{Y_{i}(1),Y_{i}(0)\}$. In the design-based framework, we implicitly condition on these observed potential outcomes, i.e., we treat $\{Y_{i}(1),Y_{i}(0)\}$ as fixed. Since $Y_{i}$, our observed response for unit $i$, is a function of its treatment assignment $W_{i}$, we can write
$Y_{i}=W_{i}Y_{i}(1)+(1-W_{i})Y_{i}(0).$
Since we implicitly condition on the potential outcomes $Y_{i}(1),Y_{i}(0)$, $Y_{i}$ is random only via the treatment assignments $W_{i}$. Let $\tau_{i}=Y_{i}(1)-Y_{i}(0)$. Then, the Average Treatment Effect (ATE) at time $t$ is:
Definition 3 (Average Treatment Effect (ATE) at time $t$).
$\bar{\tau}_{t}\coloneqq\frac{1}{t}\sum_{i=1}^{t}\tau_{i}.$
Our objective is to generate a confidence sequence for $\bar{\tau}_{t}$ in MAB experiments with guarantees on statistical power.
3.1 Extensions to Time Series Experiments
Our notation for unit $i$’s potential outcomes $\{Y_{i}(1),Y_{i}(0)\}$ implies that unit $i$’s outcome does not depend on the entire treatment assignment path. However, we only use this notation for ease of presentation. Because we adopt a design-based framework, and hence, treat the potential outcomes as fixed and non-random, our theoretical results for the MAD presented in Section 5 are agnostic to the actual outcome distribution over $i$. Hence, all of our theoretical results immediately extend to time series experiments; see Bojinov and Shephard (2019); Bojinov et al. (2023) for a formalization of the causal estimands and estimators in such settings.
4 The Mixture Adaptive Design
In this section, we formalize our proposed experimental design and state our main result.
As in the MAB literature, we assume that the treatment assignment probabilities at a given time $t$ can depend on previously observed data up to time $t-1$. Formally, let $H_{t}\coloneqq\{W_{i},Y_{i}\}_{i=1}^{t}$. For any arbitrary treatment assignment algorithm $\mathcal{A}$, even those that are adaptive and/or do not satisfy probabilistic treatment assignments, define the probability that the adaptive algorithm assigns treatment $w$ at time $i$ as $p_{i}^{\mathcal{A}}(w)\coloneqq\mathbb{P}_{\mathcal{A}}\left(W_{i}=w\mid H_{i-1}\right)$, where $\mathbb{P}_{\mathcal{A}}$ denotes probability taken with respect to $\mathcal{A}$. For instance, if the user wanted to set $\mathcal{A}$ as a Thompson sampling for their experiment, $p_{i}^{\mathcal{A}}(w)$ would be the assignment probability that Thompson sampling would assign treatment $w$ to unit $i$ given $H_{i-1}$.
Definition 4 (Mixture Adaptive Design (MAD)).
For any treatment assignment algorithm $\mathcal{A}$ and a real-valued sequence $\delta_{i}\in(0,1]$, $w\in\{0,1\}$, then the probability that the MAD will assign treatment $w$ at time $i$ is:
$p^{\text{MAD}}_{i}(w)\coloneqq\mathbb{P}_{\text{MAD}}\left(W_{i}=w\mid H_{i-1}\right)=\delta_{i}\left(\frac{1}{K}\right)+(1-\delta_{i})p_{i}^{\mathcal{A}}(w).$
Intuitively, the MAD “mixes” a (possibly adaptive and non-probabilistic) assignment algorithm $\mathcal{A}$ with a Bernoulli design. For instance, assuming $\mathcal{A}$ is a Thompson sampler, then the MAD will assign treatment $w$ according to a Bernoulli design with probability $\delta_{i}$, and according to the Thompson sampler with probability $1-\delta_{i}$, for all $i,w$. Note, $p^{\text{MAD}}_{i}(w)>0$ even for deterministic assignment algorithms like UCB because $\delta_{i}>0$ for all $i$, thus avoiding divisions by zero when computing IPW estimators. For instance, if the UCB algorithm has $p_{i}^{\mathcal{A}}(1)=0$, then $p^{\text{MAD}}_{i}(1)=\delta_{i}\left(\frac{1}{2}\right)>0$. Note, Definition 4 can easily be extended to the $K\geq 2$ treatment settings with $w\in\{0,1,...,K-1\}$ by setting $p^{\text{MAD}}_{i}(w)\coloneqq\mathbb{P}_{\text{MAD}}\left(W_{i}=w\mid H_{i-1}\right)=\delta_{i}\left(\frac{1}{K}\right)+(1-\delta_{i})p_{i}^{\mathcal{A}}(w)$.
The intuitive idea behind the MAD is that if we can balance the Bernoulli design and the adaptive design via $\delta_{i}$, we can gain the ATE precision while maintaining some of the regret minimization of the bandit algorithm. For simplicity of notation, we suppress $p^{\text{MAD}}_{i}(w)$’s dependence on $\mathcal{A}$.
We now define our ATE estimators. Let $\mathcal{F}_{t}$ be the sigma-algebra that contains all pairs of potential outcomes $\{Y_{i}(1),Y_{i}(0)\}_{i=1}^{t}$ and all observed data $\{W_{i},Y_{i}\}_{i=1}^{n}$ where $n\leq t$.
Based on the estimator for $\tau_{i}$ proposed in Bojinov and Shephard (2019); Bojinov et al. (2021) for adaptive experiment settings, we set
$\hat{\tau}_{i}\coloneqq\frac{\mathbbm{1}\{W_{i}=1\}Y_{i}}{p^{\text{MAD}}_{i}(1)}-\frac{\mathbbm{1}\{W_{i}=0\}Y_{i}}{p^{\text{MAD}}_{i}(0)},$ (1)
which is an unbiased estimator for $\tau_{i}$:
$\mathbb{E}\left[\hat{\tau}_{i}\mid\mathcal{F}_{i}\right]={\tau}_{i},$
where the expectation is taken with respect to the treatment assignment mechanism. Hence,
$\frac{1}{t}\sum_{i=1}^{t}\mathbb{E}\left[\hat{\tau}_{i}\mid\mathcal{F}_{i}\right]=\bar{\tau}_{t},$
and we define our estimator of the Average Treatment effect at time $t$ as:
$\hat{\bar{\tau}}_{t}\coloneqq\frac{1}{t}\sum_{i=1}^{t}\hat{\tau}_{i},$ (2)
which is an unbiased estimator for $\bar{\tau}_{t}$. Additionally, we have that
$Var(\hat{\tau}_{i}\mid\mathcal{F}_{i})\leq\sigma^{2}_{i}\text{, where }\sigma^{2}_{i}\coloneqq\frac{Y_{i}(1)^{2}}{p^{\text{MAD}}_{i}(1)}+\frac{Y_{i}(0)^{2}}{p^{\text{MAD}}_{i}(0)},$ (3)
and thus, a natural unbiased estimator for $\sigma^{2}_{i}$ is
$\hat{\sigma}^{2}_{i}\coloneqq\frac{Y_{i}(1)^{2}\mathbbm{1}\{W_{i}=1\}}{(p^{\text{MAD}}_{i}(1))^{2}}+\frac{Y_{i}(0)^{2}\mathbbm{1}\{W_{i}=0\}}{(p^{\text{MAD}}_{i}(0))^{2}}.$ (4)
Equations (3) and (4) follow from Bojinov and Shephard (2019); Bojinov et al. (2021), which propose the estimator of Equation (4) because the closed form does not admit a natural unbiased estimator. Let $S_{t}\coloneqq\sum_{i=1}^{t}\sigma^{2}_{i}$ and $\hat{S}_{t}\coloneqq\sum_{i=1}^{t}\hat{\sigma}^{2}_{i}$.
4.1 Extension to Batched Assignment Settings
Often, it is more practical to update the treatment assignment probabilities of an adaptive algorithm after observing a batch of units (Gao et al. 2019; Zhang et al. 2020; Nabi et al. 2022; Gong and Simchi-Levi 2023; Ren and Zhou 2024). Assume we observe a sequence of batches, where for each batch $j$, we have a (non-random and finite) batch size of $B$. So, for each batch $j$, we observe $H^{\text{batched}}_{j}\coloneqq\left\{W^{(j)}_{i},Y^{(j)}_{i}\right\}_{i=1}^{B}$, where $W^{(j)}_{i}\in\{0,1\}$ and $Y^{(j)}_{i}$ are the treatment assignment and outcome for unit $i$ in batch $j$, respectively. We assume the treatment assignment probabilities are fixed within a batch, i.e., the treatment assignment probabilities are only (adaptively) updated after observing a batch of $B$ units.
Let $p_{j}^{\mathcal{A}_{\text{batched}}}(w)\coloneqq\mathbb{P}(W^{(j)}_{i}=w\mid H^{\text{batched}}_{j-1})$ be the assignment probability to treatment $w$ for unit $i$ in the $j$th batch for any user-provided (adaptive) batched algorithm $\mathcal{A}_{\text{batched}}$. We define the Mixture Adaptive Design for Batched Assignment as:
Definition 5 (Mixture Adaptive Design for Batched Assignment).
Given a batched (adaptive) assignment algorithm $\mathcal{A}_{\text{batched}}$ and a real-valued sequence $\delta_{j}\in(0,1]$, for $w\in\mathcal{W}$, the assignment probability to treatment $w$ in batch $j$ is:
$p^{\text{MAD}_{\text{batched}}}_{j}\left(w\right)=\delta_{j}\left(\frac{1}{2}\right)+(1-\delta_{j})p_{j}^{\mathcal{A}_{\text{batched}}}(w).$
where once again, we can easily extend this definition to $K\geq 2$ treatments where we would have $p^{\text{MAD}_{\text{batched}}}_{j}\left(w\right)=\delta_{j}\left(\frac{1}{K}\right)+(1-\delta_{j})p_{j}^{\mathcal{A}_{\text{batched}}}(w)$.
Hence, in batched settings, treatment assignment probabilities update after each batch of $B$ units rather than after every individual observation. Hence, in this setting, treatments within a batch are assigned independently, i.e., for each unit $i$ in batch $j$, we assign treatment by rolling a $K$-sided dice where each treatment $w\in\mathcal{W}$ has probability $p^{\text{MAD}_{\text{batched}}}_{j}\left(w\right)$ of appearing, and $p^{\text{MAD}_{\text{batched}}}_{j}\left(w\right)$ is fixed within the batch $j$. In fact, we can think of Definition 5 as a more general version of Definition 4, where Definition 4 is just Definition 5 with the batch size $B=1$.
5 Theoretical Guarantees of the MAD
In this section, we provide both inferential and regret guarantees for the MAD. We will present results for the MAD with binary treatments in Definition 4, and fully formalize and prove the analogous results for $K\geq 2$ treatments and batched assignment settings in Appendix C and D.
5.1 Inferential Guarantees
We first state the assumptions necessary to establish our inferential results.
Assumption 1 (Bounded (Realized) Potential Outcomes).
There exists $M\in\mathbb{R}$ such that
$\limsup_{t\to\infty}|Y_{t}(w)|\leq M<\infty$
for all $w\in\mathcal{W}$.
This assumption is used in existing work on CSs for the ATE Howard et al. (2021); Ham et al. (2023) and commonly assumed in design-based causal inference settings Bojinov and Shephard (2019); Bojinov et al. (2021); Lei and Ding (2021). Note, this is an assumption on the realized potential outcomes. While this assumption may seem limiting, the realized outcomes in any real-world experiment are almost always bounded, as the limitations of computing precision guarantee that any realized outcome collected using existing computing resources will be bounded by, e.g., the highest floating point number via IEEE-754 standards. Hence, even if $Y_{t}(w)$ were drawn from a Gaussian distribution using existing computing resources, the realized potential outcomes will never exceed this upper limit on floating point precision, and therefore we can argue that this assumption is satisfied. Additionally, the confidence sequence we provide in Theorem 1 does not take in $M$ as input, and hence, the value of $M$ need not be known by the experimenter and does not affect the width of the confidence sequence.
We also require that the average conditional variance of our estimator $\frac{1}{t}\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}\mid\mathcal{F}_{i})$ is not vanishing. Specifically, we require that the cumulative conditional variances $\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}\mid\mathcal{F}_{i})$ are growing at at least a linear rate in $t$. Define $\Omega(t)$ such that if $f(t)=\Omega(t)$, there exists $k>0$ and $t_{0}>0$ such that for all $t\geq t_{0}$, $f(t)\geq kt$.
Assumption 2 (At Least Linear Rate of Cumulative Conditional Variances).
$\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}\mid\mathcal{F}_{i})=\Omega(t).$
Assumption 2 states that the sums of the conditional variances go to infinity at least as fast as a constant rate. Existing works such as Ham et al. (2023) and Waudby-Smith et al. (2023) require that $\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}\mid\mathcal{F}_{i})\to\infty$, but do not assume a specific rate. For instance, they allow for the average conditional variance to vanish superlinearly, which would violate Assumption 2. Although Assumption 2 is slightly stronger, this assumption should be satisfied in most realistic experimental settings so long as we do not have an adversarial sequence $(Y_{i}(1),Y_{i}(0))_{i=1}^{\infty}$ which would “cancels out” the rate of the $p_{i}^{\text{MAD}}(w)$’s. For instance, if there existed some time $t$ such that beyond $t$, $Y_{i}(1),Y_{i}(0)$ are constantly 00, this assumption does not hold. However, such scenarios would be unusual in practice and may indicate practical issues with the experiment. Importantly, this assumption is the only condition we impose on the user-chosen bandit algorithm (the $p^{\mathcal{A}}_{i}(w)$ in $p_{i}^{\text{MAD}}(w)$), and therefore, we expect our result to be valid for nearly any bandit algorithm.
Finally, for $a\geq 0$, let $\delta_{i}=\omega(1/i^{a})$ to mean $\delta_{i}i^{a}\to\infty$, i.e., that $1/\delta_{i}=o(i^{a})$. Intuitively, we can think of $\delta_{i}$ as a sequence approaching zero slower than $1/i^{a}$.
Theorem 1.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)$ as in Definition 4 with respect to some treatment assignment algorithm $\mathcal{A}$ and $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 2 hold. Then $(\hat{\bar{{\tau}}}_{t}\pm\hat{V}_{t})$ where
$\hat{V}_{t}=\sqrt{\frac{2(\hat{S}_{t}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}\eta^{2}+1}}{\alpha}\right)}$
is a valid $(1-\alpha)$ asymptotic CS for $\bar{\tau}_{t}$ and $\hat{V}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}0$.
The proof is provided in Appendix A. Intuitively, Theorem 1 holds because a Bernoulli Design provides this guarantee (simply set $\delta_{i}=1$ above), so a design that stochastically injects a Bernoulli design into a MAB experiment will maintain the same guarantee as long as the experiment does not deviate from the Bernoulli design too quickly, i.e., the rate at which $\delta_{i}$ approaches 00 is controlled.
At a high level, proving the validity of Theorem 1 proceeds by showing that $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$ along with Assumptions 1 and 2 ensure that the ATE estimator of Equation (1) satisfies a Lindeberg-type uniform integrability (see Lemma A.1 in Appendix A), thus allowing us to apply the universal asymptotic CS of Waudby-Smith et al. (2023) in this setting. The shrinking variance result proceeds by showing that $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$ and Assumption 1 guarantees that $\hat{S}_{t}\log\hat{S}_{t}=o(t^{2})$ almost surely, and hence, $\hat{V}_{t}$ is shrinking towards 00 almost surely. Note, this CS is not a function of $M$ in Assumption 1; $M$ is only used to ensure the Lindeberg-type uniform integrability result and to control the asymptotic rate of $\hat{S}_{t}$. Therefore, $M$ can be set arbitrarily large to satisfy Assumption 1 without affecting the width of our CS.
Practically, Theorem 1 enables experimenters to perform bandit experiments using the MAD and, as long as Assumptions 1 and 2 hold, be guaranteed both valid and powerful inference. While Ham et al. (2023) and Howard et al. (2021) propose CSs for the ATE which allow for adaptive assignment, their CS requires probabilistic assignment assumption for their validity guarantees. Both Ham et al. (2023) and Howard et al. (2021) also require Assumption 1, so Theorem 1 removes the probabilistic assignment assumption entirely while only requiring a slightly stronger assumption on the cumulative conditional variances (which should be easily satisfied in most practical settings, as discussed above). Additionally, we prove the novel result that the width of our asymptotic CS is shrinking by deriving the rate of $\hat{S}_{t}$ for the MAD. While Ham et al. (2023) discusses how their asymptotic CS scales with the variance of the estimator and state that they expect it to shrink in many adaptive settings, they do not prove that $\hat{S}_{t}$ has the correct rate to ensure a shrinking CS, even with probabilistic assignments.
As discussed in Ham et al. (2023), experimenters at companies such as Netflix often use the first time zero is outside of a CS as a stopping rule for their online experiments. We show that as long as $\bar{\tau}_{t}$ is truly non-zero in the long term, such a stopping rule will occur in finite time.
Let $T_{MAD}\coloneqq\inf_{t}\left\{t:0\notin(\hat{\bar{\tau}}_{t}\pm\hat{V}_{t})\right\}$, i.e., this is the first time 00 is not within the confidence sequence specified in Theorem 1.
Theorem 2.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)$ as in Definition 4 with respect to some treatment assignment algorithm $\mathcal{A}$ and $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 2 hold. Assume $\bar{\tau}_{t}\to c$ for some $|c|>0$ as $i\to\infty$. Then,
$\mathbb{P}\left(T_{MAD}<\infty\right)=1.$
The proof is provided in Appendix B. The assumption that $\bar{\tau}_{t}\to c$ is satisfied immediately in stationary outcome settings, e.g., if we assume $\tau_{i}=c$ for all $i$ or if $\tau_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{D}$ for some distribution $\mathcal{D}$ with finite expectation. In non-stationary settings, this assumption intuitively says that the non-stationarity should be limited (e.g., at the beginning of the experiment) such that the long-term behavior of $\bar{\tau}_{t}$ stabilizes towards some value.
As part of the proof for Theorem 2, we show that the CS of Theorem 1 is shrinking around the true ATE, so if $\bar{\tau}_{t}\to c$ and $|c|>0$, the CS will at some point exclude zero. Note, existing work such as Ham et al. (2023) have no coverage guarantees, not even asymptotic ones, in general bandit settings and in practice, can severely undercover if naively applied to a bandit setting, i.e., the percentage of times that the true ATE is included in their CS is far below $1-\alpha$.
5.1.1 Inference for $K\geq 2$ Treatment and Batched Assignment Settings
For any pair of treatments $w,w^{\prime}$ where a non-zero treatment effect exists, analogous results of Theorem 1 and Theorem 2 hold for the MAD with $\geq 2$ treatments for both the standard and batched assignment (Defintion 5) setting; full statement and proof are in Appendix C and D.
In particular, settings where $K>2$ showcases the advantages of our anytime valid guarantees, as the experimenter can iteratively exclude “unpromising” treatments from consideration after observing sufficient inferential evidence, thus reducing harm to the experimental units. For example, assume the experimenter has $K-1$ total treatments, $w_{1},...,w_{K-1}$ under consideration and one control, $w_{0}$. At each time step $t$, the experimenter can compute the CS of Theorem 1 for each of the $K-1$ treatment-control pairs $(w_{j},w_{0})$, $j=1,..,.K-1$. Then, if the experimenter observes sufficient evidence of strong treatment effects between certain treatments and the control (e.g., the confidence sequence excludes 00 and/or seems to be centered around a non-zero value) while others seem to have weak or possibly no treatment effects, the experimenter can decide to remove these “weak” treatments from consideration and continue the experiment with only the most promising treatments.
5.1.2 Connections to Exact Anytime Valid Inference
Since $T_{MAD}$ is a stopping time based on an asymptotic CS, a reasonable concern is that the CS at any finite time point $t$ will not have finite-sample error control, and hence, any inferential conclusion drawn based on the level-$\alpha$ asymptotic CS at $T_{MAD}$ may not be statistically valid. However, recalling the definition of an asymptotic CS in Definition 2, there is some underlying exact (i.e., non-asymptotic) CS that the asymptotic CS of Theorem 1 is converging towards almost surely. We show that, under the same assumptions of Theorem 2, this unknown, exact CS is also guaranteed to have a finite stopping time. This results implies that, if one had access to the true underlying exact CS, it would make the same inferential conclusion as the asymptotic CS, just at a later time point in the experiment.
Formally, define $T^{*}_{MAD}\coloneqq\inf_{t}\left\{t:0\notin(\hat{\bar{\tau}}_{t}\pm V^{*}_{t})\right\}$ as the analogous stopping time to $T_{MAD}$ for the exact anytime valid CS underlying the asymptotic CS of Theorem 1.
Theorem 3.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)$ as in Definition 4, $w\in\{0,1\}$, and $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 2 hold. Assume $\bar{\tau}_{t}\to c$ for some $|c|>0$ as $t\to\infty$. Then, $\mathbb{P}\left(T^{*}_{MAD}<\infty\right)=1$.
Hence, though the asymptotic CS cannot guarantee exact error control at the stopping time, it will make the same inferential conclusion as its underlying exact counterpart. In particular, we show in the proof of Theorem 3 that $V^{*}_{t}=\hat{V}_{t}+o(1)$ almost surely for $\delta_{i}=\omega(1/i^{1/4})$, and hence, the asymptotic CS approaches its exact counterpart at a constant rate which does not depend on $t$. In practice, the user can simply run the experiment for a few additional samples after $T_{MAD}$ is observed to check whether the the CSs for $t>T_{MAD}$ still excludes zero.
However, we can also use the MAD to perform exact anytime valid inference. We prove the same stopping time guarantee for the MAD using the exact anytime valid CS for ATE estimation provided in Corollary 2 of Howard et al. (2021).
Define $\tilde{T}_{MAD}\coloneqq\inf_{t}\left\{t:0\notin(\hat{\bar{\tau}}_{t}\pm u(W_{t})/t\right\}$ where $u(W_{t})$ is defined as on Corollary 2 of Howard et al. (2021) and $W_{t}=\sum_{i=1}^{t}(\hat{\tau}_{i}-\hat{\bar{\tau}}_{i})^{2}$.
Lemma 3.1.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)$ as in Defintion 4 with $w\in\{0,1\}$, where there exists some $10<p_{min}<$ 1 such that $p_{i}^{\text{MAD}}(w)\geq\min\{p_{min},1-p_{min}\}$ for all $i\in[1,\infty)$. Assume Assumptions 1 and 2 hold. Assume $\bar{\tau}_{t}\to c$ for some $|c|>0$ as $t\to\infty$. Then, $\mathbb{P}\left(\tilde{T}_{MAD}<\infty\right)=1$.
However, using the exact CS has a number of limitations in practice. Firstly, to guarantee exact anytime validity and shrinking width, this exact CS requires the MAD to have probabilistic treatment assignments, i.e., that all treatment assignment probabilities are bounded within some $[p_{min},1-p_{min}]$. This probabilistic treatment assignment condition greatly restricts the user’s choices of $\delta_{i}$ and the assignment algorithm $\mathcal{A}$, since the treatment assignment probabilities can no longer approach 00 or $1$ as more samples are observed. The only way to ensure $p_{i}^{\text{MAD}}(w)\geq\min\{p_{min},1-p_{min}\}$ is to either choose a $\delta_{i}$ which can approach 00 but an $\mathcal{A}$ which has assignment probabilities bounded away from 00 and $1$, which would exclude most common adaptive algorithms like Thompson sampling and UCB, or to use any $\mathcal{A}$ but restrict $\delta_{i}\geq\min\{p_{min},1-p_{min}\}$, i.e., MAD is forced to assign treatment via a Bernoulli design at a greater than $p_{min}$ rate for all time $t$.
Secondly, as discussed in previous works (Ham et al. 2023; Waudby-Smith et al. 2023), exact CSs such as those of Howard et al. (2021) not only require restrictive distributional assumptions on the data, such as known bounds on the estimator and/or its moments, to establish validity, but can also generate far less precise inference than asymptotic CSs. In fact, we show in Appendix G that the CS of Howard et al. (2021) shrinks towards the true ATE far slower than our asymptotic CS of Theorem 1 and thus, requires far more samples to produce precise inference. Importantly, our asymptotic CS tends to maintain the desired coverage and error control even in moderate sample sizes. In fact, we find in Sections 7 and G that the coverage and error control of our asymptotic CS is nearly indistinguishable from that of a Bernoulli design across a wide range of experimental settings, even for relatively small sample sizes given the scale of most online experiments ($<100$).
The statistics literature has a long history of using asymptotic tests such as the Wald test in finite-sample settings Van der Vaart (2000), where the primary and largely accepted justification used is that such asymptotic tests can well approximate finite-sample settings once a sufficient sample size has been reached, and thus, will provide close-to-exact error control. We show through our empirical results that our asymptotic CS provides essentially exact error control even for small sample sizes. Based on these results, and the inferential congruence guarantee of Theorem 3, we believe it is statistically justifiable to use of the asymptotic CS of Theorem 1 in practice, especially given the scale of most online experiments.
5.2 Regret Guarantees
In many applications, experimenters run adaptive algorithms with the goal of simultaneously minimizing harm and conducting inference, usually with some prioritization between these two goals. Because the MAD incorporates a Bernoulli design into an adaptive assignment algorithm to gain inferential guarantees, a natural next question is: what is lost in terms of regret in order to gain this inference? We show that the regret of the MAD can be represented as a weighted sum between the regret of the underlying adaptive algorithm $\mathcal{A}$ and the regret of a Bernoulli design, where the weights are given by $\delta_{i}$.
Assume binary treatments with the control arm having mean outcome $\mu_{0}$ and the treatment arm having mean outcome $\mu^{*}$ with $\mu^{*}\geq\mu_{0}$. Let $Y_{i}^{\text{obs}}$ be the observed outcome/reward at time $t$. Let $\mathbb{E}_{\mathcal{A}}[Y_{i}^{\text{obs}}]$, $\mathbb{E}_{\text{Bern}}[Y_{i}^{\text{obs}}]$, and $\mathbb{E}_{\text{MAD}}[Y_{i}^{\text{obs}}]$ denote the expected outcome (often referred to as the reward in the reinforcement learning literature) taken with respect to the an adaptive assignment algorithm $\mathcal{A}$, a Bernoulli design, and the MAD using $\mathcal{A}$ as the adaptive algorithm, respectively. Define the regret of a design $\pi$ (often referred to as the policy in the reinforcement learning literature) at time $i$ as $R_{i}(\pi)=\mu^{*}-\mathbb{E}_{\pi}[Y_{i}^{\text{obs}}]$, where $\pi\in\{\mathcal{A},\text{Bern},\text{MAD}\}$. Hence, the regret of a design $\pi$ is the expected difference in outcome between the outcome-maximizing design (which assigns treatment with probability $1$) and $\pi$. Then,
$R_{i}(\text{MAD})$ $=\mu^{*}-\mathbb{E}_{\text{MAD}}[Y_{i}^{\text{obs}}],$
$=\mu^{*}-\left(\delta_{i}\mathbb{E}_{\text{Bern}}[Y_{i}^{\text{obs}}]+(1-\delta_{i})\mathbb{E}_{\mathcal{A}}[Y_{i}^{\text{obs}}]\right),$
$=\mu^{*}-\left(\delta_{i}\left(\frac{\mu_{0}+\mu^{*}}{2}\right)+(1-\delta_{i})\mathbb{E}_{\mathcal{A}}[Y_{i}^{\text{obs}}]\right),$
and hence,
$R_{i}(\text{MAD})-R_{i}(\mathcal{A})$ $=\mu^{*}-\left(\delta_{i}\left(\frac{\mu_{0}+\mu^{*}}{2}\right)+(1-\delta_{i})\mathbb{E}_{\mathcal{A}}\left[Y_{i}^{\text{obs}}\right]\right)-\left(\mu^{*}-\mathbb{E}_{\mathcal{A}}\left[Y_{i}^{\text{obs}}\right]\right)$
$=\delta_{i}\left(\mathbb{E}_{\mathcal{A}}\left[Y_{i}^{\text{obs}}\right]-\frac{\mu_{0}+\mu^{*}}{2}\right)$
Hence, at each time point, the difference in regret between the MAD and the adaptive algorithm $\mathcal{A}$ shrinks towards 00 at a $\delta_{i}$ rate. In particular, the above results implies that the MAD and the adaptive algorithm will asymptotically make the same decision at each time point, so if the adaptive algorithm correctly identifies the best arm, so too will the MAD.
In contrast with experimental design works such as Kasy and Sautmann (2021) which do not allow the user to tune the degree of regret minimization in the algorithm and Simchi-Levi and Wang (2023), where this tuning is a function of various hyperparameters which may tradeoff with one another, the regret of the MAD is explicitly controlled by a single user-chosen parameter, $\delta_{i}$.  [^1] Recall also that Kasy and Sautmann (2021) and Simchi-Levi and Wang (2023) each present a specific design for a specific adaptive algorithm, and thus, cannot be used generally for any adaptive algorithm. Additionally, both these works only characterize the asymptotic behavior of their design’s regret. Using the MAD, the user can exactly quantify at each time point the loss in regret between the MAD and its underlying adaptive algorithm through $\delta_{i}$ and thus, has complete control over the rate at which that loss shrinks over time. To our knowledge, no other experimental design for adaptive experiments provides such a finite-sample characterization of regret.
6 Recommendations for Setting $\delta_{i}$
Theorems 1 and 2 hold for any $\delta_{i}=\omega(\frac{1}{i^{1/4}})$, therefore, the MAD provides flexibility to select $\delta_{i}$ based on the specifics of the problem. In this section, we provide a few recommendations of $\delta_{i}$ based on the experimenter’s priorities.
Example 1: $\delta_{i}=1/i^{0.24}$; such a $\delta_{i}$ would allow the experimenter to aggressively prioritize the bandit design (and hence regret minimization) while still maintaining validity and shrinking width guarantees.
Example 2: $\delta_{i}=c$ for some $c\in(0,1]$; such a $\delta_{i}$ would maintain a constant rate of priority in the Bernoulli design. Such a $\delta_{i}$ can be valuable in settings where the experimenter may want to stop the experiment as soon as possible (e.g., due to an external deadline, limited or expensive resources, etc.) via some stopping rule (e.g. when 00 is outside the interval), and so, would like to generate precise inferential results as quickly as possible while still allowing for some regret minimization. Note that choosing $c=1$ reduces to a Bernoulli design.
Example 3: $\delta_{i}=\max\{\tilde{\delta}_{i},c\}$ for some $c\in(0,1]$, $\tilde{\delta}_{i}$ a sequence in $(0,1]$; such a $\delta_{i}$ would prioritize the bandit design up until a point, then maintain a constant rate of priority on the Bernoulli design in the long run, thus never fully prioritizing the bandit over the Bernoulli design. For instance, since experimenters often have an upper bound $N$ in mind for the maximum time they would ideally like to run the experiment (e.g., due to an external deadline, limited or expensive resources, etc.), they could set $\tilde{\delta}_{i}=i^{0.24}$ and $c=1/N^{0.24}$. Such a $\delta_{i}$ can be valuable in settings where the experimenter primarily prioritizes regret minimization, and hence, initially chooses an aggressive $\tilde{\delta}_{i}$ to see if they can achieve statistically significant results before the $N$ samples/time steps. If such results are not achieved by $N$ samples, then the experimenter will maintain a constant rate of Bernoulli assignment from that point onward to sharpen the inference further.
Thus, experimenters can flexibly determine a $\delta_{i}$ that balances the trade-off between regret minimization and statistical power throughout the experiment. Note also, setting $\delta_{i}$ as in Example 2 and 3 is similar to clipping in which the assignment probabilities of an stochastic adaptive assignment algorithm like TS are clipped to never fall below a certain value $p_{i}>0$ Jin et al. (2020); Dimakopoulou et al. (2021). For instance, if we set $\delta_{i}=c$ for some $c>0$, then for a two-armed MAB, TS would choose, at each timepoint $i$, arm $1$ with probability $(1-c)p^{\mathcal{A}}_{i}(1)+c/2$ and arm 00 with probability $1$ with probability $(1-c)p^{\mathcal{A}}_{i}(0)+c/2$, and hence, the assignment probability is forced to never drop below $c/2$, which is similar to clipping setting $p_{i}=c/2$. Similarly, setting $\delta_{i}=\max\{\tilde{\delta}_{i},c\}$ would be like clipping with a $p_{i}$ that varies with $i$ but is still lower bounded by $c/2$.
Recall from Section 1 that most existing works in bandit experimental design prescribe specific algorithms that do not allow for user tuning at all Kasy and Sautmann (2021) or limitations on when during the experiment the user can tune the exploration Simchi-Levi and Wang (2023). Although potentially useful as out-of-the-box algorithms, such approaches provide the user less flexibility to tune the algorithm based on their specific priorities or experimental constraints.
7 Simulation Study
Empirically, we find that the CS of Theorem 1 using the MAD with an appropriately set $\delta_{i}$ maintains finite-sample anytime validity while powerfully detecting non-zero ATEs. We first consider a two-arm bandit with $Y_{i}(1)\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Bern(p_{1})$ and $Y_{i}(0)\stackrel{{\scriptstyle i.i.d.}}{{\sim}}Bern(p_{0})$, so for all $t$, $\bar{\tau}_{t}=\bar{\tau}=p_{1}-p_{0}$. simply We implement both UCB and a Thompson sampler with uniform priors on both arms. For the MAD, we consider the MAD with $\delta_{i}=\frac{1}{i^{0.24}}$, which we call the unclipped MAD, and with $\delta_{i}=\max(\frac{1}{i^{0.24}},0.2)$ like in Example 2 of Section 6, which we call the clipped MAD. We also consider two baselines: a Bernoulli design (i.e., $\delta_{i}=1$) and a Standard Bandit design (i.e., $\delta_{i}=0$). Note, we can only implement the Standard Bandit design with the Thompson sampler since UCB is a deterministic algorithm, i.e., it generates treatment assignment probabilities that are either 00 or $1$ at each time step, and hence, the IPW estimator cannot be computed due to the division by zero.
At each time step, the adaptive algorithm (either UCB or the Thompson sampler) pulls a single arm and observes the outcome. For each applicable design, the CS of Theorem 1 is computed over $T=10,000$ time steps and $N=100$ independent replicates/random seeds. Note that since we do not adjust the assignment probabilities for the Standard Bandit Design, there are no validity guarantees for this design. We set $\alpha=0.05$ for computing the CS throughout. For the $\eta$ parameter of the CS of Theorem 1, we set $\eta=\sqrt{\frac{-2\log(\alpha)+\log(-2\log(\alpha)+1)}{t^{*}}}\approx 0.028$ with $t^{*}=10{,}000$; we choose this value following the recommendations in Appendix B.2 of Waudby-Smith et al. (2023) on setting $\eta$ to optimize the width of the CS for a specific time $t^{*}$, where we take $t^{*}$ to be the end of the time horizon we set for this experiment, $T$. We perform this analysis for $(p_{0},p_{1})=\{(0.5,0.5),(0.6,0.8),(0.2,0.8)\}$ so the ATE = $0,0.2,0.6$ respectively.
As shown in Figure 1, the width of the CSs under both the clipped and unclipped MADs shrinks much faster than the Standard Bandit Design, nearly matching that of the Bernoulli design. Figure 2 depicts the empirical coverage at each time point $t$, which we calculate as the proportion of CSs up to and including $t$ for which the true ATE is inside the CS at each time point, i.e., we compute $\frac{1}{t}\sum_{j=1}^{t}\mathbbm{1}\left\{\bar{\tau}\in\left(\hat{\bar{\tau}}_{j}\pm\hat{V}_{j}\right)\right\}$. We use this empirical coverage to estimate our anytime valid error control as given in Definition 1. The clipped and unclipped MADs have $1-\alpha$ error control across all time as desired, even when $t$ is small, showing that the CS of Theorem 1 can maintain proper finite-sample anytime validity even though it is an asymptotic CS. Notably, the Standard Bandit design does not maintain the proper coverage when the ATE is non-zero, exhibiting how naively using a Standard Bandit design cannot ensure anytime validity. When the ATE is zero, the Standard Bandit design does maintain the proper coverage, intuitively because the bandit algorithm never strongly favors one arm over the other since the two arms are indistinguishable.
Hence, the Standard Bandit design tends to produce wider CSs than the MAD but still lacks the correct coverage when the ATE is non-zero. Intuitively, this result occurs because as time increases, the assignment probability for arm 00 becomes negligibly small so that the algorithm essentially only draws from arm $1$. Hence, the ATE estimator $\hat{\bar{\tau}}_{t}$ as in Equation (2) is computed essentially only using observations from arm $1$. This skewing is worse when the ATE is larger because the assignment probability for arm 00 more quickly approaches 00.
Recall, an experimenter may define a stopping rule such that she stops at the first time 00 is not in the CS. Figure 3 shows the proportion of replicates which do not contain 00 at time $t$. The MAD excludes 00 with much fewer samples than the Standard Bandit design for both choices of $\delta_{t}$, essentially matching that of the Bernoulli design when the ATE is large. Notably, when the ATE is non-zero, a significant proportion of the CSs for the Standard Bandit design never exclude 00 by the end of the time horizon.
Finally, Figure 4 shows the time-average reward of each of the designs, which we compute as $\frac{1}{t}\sum_{i=1}^{t}Y_{t}^{\text{obs}}$ for each $t$. As expected based on the results of Section 5.2, the reward of the MAD quickly approaches that of the Standard Bandit Design. In conclusion, our results show that the MAD maintains anytime validity even in finite samples, and has a CS width and stopping time behavior that is similar to the Bernoulli design while achieving reward that is closer to the Standard Bandit design. Note also that, as shown in Figure 1 and Figure 4, the clipped MAD can have a slightly smaller width over time but also a slightly smaller reward than the unclipped MAD. This makes sense as the clipped MAD enforces greater exploration rather than exploitation. We expect that with a more aggressive clipping rate, this difference will be more stark. The experimenter can choose $\delta_{t}$ which balances this tradeoff given their problem specifics.
Figure 1: Average value of $\hat{V}_{t}$, which defines the width of the CS of Theorem 1, across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Bernoulli outcome model, as described in Section 7. Error bands depict ±$\pm 2$ SEs.
Figure 2: Empirical coverage of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Bernoulli outcome model, as described in Section 7. $1-\alpha$ is depicted by the dotted black line. Error bands depict ±$\pm 2$ SEs.
Figure 3: The proportion of replicates which do not contain 00 at time $t$ of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Bernoulli outcome model, as described in Section 7. $1-\alpha$ is depicted by the dotted black line. Note, when the ATE $=0$, we are simply measuring the empirical coverage, which we want to control at $1-\alpha$. Error bands depict ±$\pm 2$ SEs.
Figure 4: The time averaged reward of each of the designs across $N=100$ random seeds under a two-armed bandit setting with a Bernoulli outcome model, as described in Section 7. Error bands depict ±$\pm 2$ SEs.
We also perform the analogous experiments under Normal, Cauchy, and t-distributed outcome models; see Appendix G. Similar to the Bernoulli outcome results, the MAD generally maintains finite-sample validity and can be robust even in misspecified settings where the true outcome model is t-distributed or even Cauchy and we use a Thompson sampler assuming a Normal outcome model. In contrast, the Standard Bandit design generally does not have finite-sample validity, even in well-specified settings.
Using the clipped MAD, we also perform a subset of the Bernoulli outcome experiments using the non-asymptotic CS of Howard et al. (2021); see Appendix G. As expected, the CS of Howard et al. (2021) is generally much wider than our asymptotic CS and tends to overcover. As our asymptotic CS maintains finite-sample anytime validity across a wide range of outcomes models and sample sizes and has a smaller width, we recommend the use of our asymptotic CS in practice.
The script used to reproduce all simulation results is provided at https://github.com/biyonka/mixture_adaptive_design.
8 Conclusion
We propose the Mixture Adaptive Design (MAD), an experimental design for valid and powerful anytime-valid, design-based causal inference in adaptive experiments. By controlling the rate of a bandit algorithm’s adaptive assignment probabilities via “mixing” with a Bernoulli design, the MAD provides a confidence sequence for the ATE that is asymptotically valid with width guaranteed to shrink to zero for nearly any choice of bandit algorithm. As existing approaches do not guarantee validity in general bandit settings, the MAD provides a practical tool for online and anytime valid causal inference in both standard and batched bandit settings.
9 Acknowledgements
We thank Nathan Cheng for his helpful discussions and feedback.
SUPPLEMENTARY MATERIAL
Appendix A Proof of Theorem 1
To prove Theorem 1, we must first show that $(\hat{\tau}_{i})_{i=1}^{\infty}$ satisfies a Lindeberg-type uniform integrability condition (Condition L-2 of Waudby-Smith et al. (2023)).
Lemma A.1.
Let $(\hat{\tau}_{t})_{t=1}^{\infty}$ be a sequence of random variables where $W_{t}=w$ with probability $p_{t}^{\text{MAD}}(w)$ for $w\in\{0,1\}$ and $\delta_{t}\in(0,1]$ such that $\delta_{t}=\omega\left(\frac{1}{t^{1/4}}\right)$. Assume Assumptions 1 and 2 hold. Then, $(\hat{\tau}_{t})_{i=1}^{\infty}$ satisfies a Lindeberg-type uniform integrability condition, i.e., there exists $\kappa\in(0,1)$ such that
[…]
where $B_{t}=\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}\mid\mathcal{F}_{i})$.
Proof.
By Assumption 1,
$\left|\hat{{\tau}}_{t}\right|=\left|\frac{Y_{t}(1)\mathbbm{1}\{W_{t}=1\}}{p^{\text{MAD}}_{t}(1)}-\frac{Y_{t}(0)\mathbbm{1}\{W_{t}=0\}}{p^{\text{MAD}}_{t}(0)}\right|\leq\frac{2M}{\min(p^{\text{MAD}}_{t}(0),p^{\text{MAD}}_{t}(1))}$
and $|{\tau}_{t}|\leq 2M$ for all $t$.
Hence,
$(\hat{{\tau}}_{t}-{\tau}_{t})^{2}$ $\leq\left(\frac{2M}{\min(p^{\text{MAD}}_{t}(0),p^{\text{MAD}}_{t}(1))}\right)^{2}+(2M)^{2}+2\left(\frac{2M}{\min(p^{\text{MAD}}_{t}(0),p^{\text{MAD}}_{t}(1))}\right)2M$
$\leq\frac{24M^{2}}{(\min(p^{\text{MAD}}_{t}(0),p^{\text{MAD}}_{t}(1)))^{2}}$
First, note that for all $w$, $p^{\text{MAD}}_{t}(w)\geq\delta_{t}(1/2)$ and so $\frac{1}{p^{\text{MAD}}_{t}(w)}=o(t^{1/4})$. So, $(\hat{{\tau}}_{t}-{\tau}_{t})^{2}=o(t^{1/2}$) almost surely.
By Assumption 2, $B_{t}=\Omega(t)$. Therefore, $B_{t}^{\kappa}=\Omega(t^{\kappa})$. Set $\kappa>1/2$. Then, there exists some $\tilde{t}$ such that for all $t\geq\tilde{t}$, $B_{t}^{\kappa}>(\hat{{\tau}}_{t}-{\tau}_{t})^{2}$ a.s.. Hence, for all $t\geq\tilde{t}$, $\mathbbm{1}\{(\hat{{\tau}}_{t}-{\tau}_{t})^{2}>(B_{t})^{\kappa}\}=0$, and so,
[…]
Theorem 1.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)$ as in Definition 4 with $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 2 hold. Then $(\hat{\bar{{\tau}}}_{t}\pm\hat{V}_{t})$ where
$\hat{V}_{t}=\sqrt{\frac{2(\hat{S}_{t}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}\eta^{2}+1}}{\alpha}\right)}$
is a valid $(1-\alpha)$ asymptotic CS for $\bar{\tau}_{t}$ and $\hat{V}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}0$.
Proof.
By Assumptions 1 and 2 imply that Lemma A.1 holds. Lemma A.1 and Assumption 2 satisfy Conditions L-1 and L-2 of Theorem 2.5 in Waudby-Smith et al. (2023), so, by Steps 1 and 2 of the proof of Theorem 2.5 in Waudby-Smith et al. (2023),
$(\hat{\bar{{\tau}}}_{t}\pm V^{*}_{t})$ where
$V^{*}_{t}=\sqrt{\frac{2(B_{t}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{B_{t}\eta^{2}+1}}{\alpha}\right)}$
is a valid $(1-\alpha)$asymptotic CS and $B_{t}=\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}\mid\mathcal{F}_{i})$.
As noted in Equation (3),
$Var(\hat{\tau}_{i}\mid\mathcal{F}_{i})\leq\sigma^{2}_{i}\text{, where }\sigma^{2}_{i}\coloneqq\frac{Y_{i}(1)^{2}}{p_{i}(1)}+\frac{Y_{i}(0)^{2}}{p_{i}(0)}.$ (5)
Hence, $(\hat{\bar{{\tau}}}_{t}\pm\tilde{V}_{t})$ where
$\tilde{V}_{t}=\sqrt{\frac{2(S_{t}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{S_{t}\eta^{2}+1}}{\alpha}\right)}$
is still a valid $(1-\alpha)$asymptotic CS where recall, $S_{t}=\sum_{i=1}^{t}\sigma_{i}^{2}$. This holds because replacing $B_{t}$ with $S_{t}$ only makes the CS wider since $S_{t}\geq B_{t}$.
As noted in Equation (4), an unbiased estimator for $\sigma^{2}_{i}$ is
$\hat{\sigma}^{2}_{i}\coloneqq\frac{Y_{i}(1)^{2}\mathbbm{1}\{W_{i}=1\}}{p^{2}_{i}(1)}+\frac{Y_{i}(0)^{2}\mathbbm{1}\{W_{i}=0\}}{p^{2}_{i}(0)}.$ (6)
and we define $\hat{S}_{t}=\sum_{i=1}^{t}\hat{\sigma}_{i}^{2}$.
To establish the validity of the CS in Theorem 1, we must show that $\frac{1}{t}\hat{S}_{t}-\frac{1}{t}S_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}0$. Then, Condition L-3 of Theorem 2.5 of Waudby-Smith et al. (2023) is satisfied and we can conclude that $(\hat{\bar{{\tau}}}_{t}\pm\hat{V}_{t})$ where
$\hat{V}_{t}=\sqrt{\frac{2(\hat{S}_{t}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}\eta^{2}+1}}{\alpha}\right)}$
is still a valid $(1-\alpha)$ asymptotic CS.
First, note that
$(\hat{\sigma}_{i}^{2})^{2}$ $\leq\left(\frac{M^{2}}{p_{i}^{MAD}(1)^{2}}+\frac{M^{2}}{p_{i}^{MAD}(0)^{2}}\right)^{2}$
$=\frac{M^{4}}{p_{i}^{MAD}(1)^{4}}+\frac{M^{4}}{p_{i}^{MAD}(0)^{4}}+2\frac{M^{2}}{p_{i}^{MAD}(1)^{2}p_{i}^{MAD}(0)^{2}}$
$\leq\frac{M^{4}}{(\delta_{i}/2)^{4}}+\frac{M^{4}}{(\delta_{i}/2)^{4}}+2\frac{M^{2}}{(\delta_{i}/2)^{4}}$
$=M^{4}\left(2^{4}\frac{1}{\delta_{i}^{4}}+2^{4}\frac{1}{\delta_{i}^{4}}+2^{5}\frac{1}{\delta_{i}^{4}}\right)$
$=o(i)$
where the last line follows because $\frac{1}{\delta_{i}}=o\left(i^{1/4}\right)$. Define $X_{i}=\hat{\sigma}_{i}^{2}-\sigma_{i}^{2}$ and note that $X_{i}$ is a martingale difference sequence. Hence,
$\mathbb{E}[X_{i}]$ $=E[(\hat{\sigma}_{i}^{2})^{2}]-(\sigma_{i}^{2})^{2}$
$\leq E[(\hat{\sigma}_{i}^{2})^{2}]$
$=o(i)$
So, $\frac{\mathbb{E}[X_{i}^{2}]}{i^{2}}=\frac{o(i)}{i^{2}}$ and hence, $\sum_{i=1}^{\infty}\frac{\mathbb{E}[X_{i}^{2}]}{i^{2}}<\infty$. For instance, if $\delta_{i}=\frac{1}{i^{a}}$ for some $40\leq a<1/$ 4, then there exists a constant $C<\infty$ and $i_{0}>0$ such that for all $i>i_{0}$, $\frac{\mathbb{E}[X_{i}^{2}]}{i^{2}}\leq Ci^{4a-2}$, and since $14a-2<-$ 1, $\sum_{i=1}^{\infty}\frac{\mathbb{E}[X_{i}^{2}]}{i^{2}}\leq C\sum_{i=1}^{\infty}i^{4a-2}<\infty$ by the p-series test.
Then, by the SLLN for martingale difference sequences (Theorem 1 of Csörgő (1968)), we can conclude that
$\frac{1}{t}\sum_{i=1}^{t}X_{i}\stackrel{{\scriptstyle a.s.}}{{\to}}0.$
Hence, Condition L-3 of Waudby-Smith et al. (2023) is satisfied and by Step 3 of the proof for Theorem 2.5 of Waudby-Smith et al. (2023), we can conclude that $(\hat{\bar{{\tau}}}_{t}\pm\hat{V}_{t})$ where
$\hat{V}_{t}=\sqrt{\frac{2(\hat{S}_{t}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}\eta^{2}+1}}{\alpha}\right)}$
is still a valid $(1-\alpha)$ asymptotic CS.
To show $\hat{V}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}0$, note that
$\hat{\sigma}_{i}^{2}$ $\leq\frac{M^{2}}{p_{i}^{MAD}(1)^{2}}+\frac{M^{2}}{p_{i}^{MAD}(0)^{2}}$
$\leq\frac{M^{2}}{(\delta_{i}/2)^{2}}+\frac{M^{2}}{(\delta_{i}/2)^{2}}$
$=\frac{4M^{2}}{\delta_{i}^{2}}+\frac{4M^{2}}{\delta_{i}^{2}}$
$=8M^{2}\frac{1}{\delta_{i}^{2}}$
$=8M^{2}o(i^{1/2})$
So, $\hat{S}_{t}\leq 8M^{2}\sum_{i=1}^{t}o(i^{1/2})$ a.s..
So, there exists positive real numbers $N$ and $x_{0}$ such that for all $t\geq x_{0}$, $\hat{S}_{t}\leq N\sum_{i=1}^{t}i^{1/2}<N\sum_{i=1}^{t}t^{1/2}=Nt^{3/2}$, and hence, we can conclude that $\hat{S}_{t}=O(t^{\frac{3}{2}})$.
We can show that $\log\left(\hat{S}_{t}\right)=o(t^{1/2})$ a.s.. For $t\geq x_{0}$,
$\frac{\log(\hat{S}_{t})}{t^{1/2}}\leq\frac{\log(Nt^{3/2})}{t^{1/2}}=\frac{\log(N)+(3/2)\log(t))}{t^{1/2}}$
and $\frac{\log(N)+(3/2)\log(t))}{t^{1/2}}\to 0$ as $t\to\infty$. Therefore, $\log(\hat{S}_{t})=o(t^{1/2})$ a.s. and we conclude that
[…]
Hence, $\hat{V}_{t}=\frac{2(\hat{S}_{t}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}\eta^{2}+1}}{\alpha}\right)=o(1)$ a.s..
Note, if $1/\delta_{i}$ was increasing at a rate faster than $i^{1/4}$ asymptotically, we are not guaranteed that $\hat{V}_{t}=o(1)$ a.s..
Appendix B Proof of Theorem 2
Theorem 2.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)$, and $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 2 hold. Assume $\bar{\tau}_{t}\to c$ for some $|c|>0$ as $i\to\infty$. Then,
$\mathbb{P}\left(T_{MAD}<\infty\right)=1.$
Proof.
We will first show that $\frac{1}{t}\sum_{i=1}^{t}\hat{\tau}_{i}\stackrel{{\scriptstyle a.s.}}{{\to}}c$. Let $u_{i}\coloneqq\hat{\tau}_{i}-\tau_{i}$.
Note that
$\mathbb{E}[u_{i}^{2}]$ $=\mathbb{E}[\hat{\tau}_{i}^{2}]-\tau_{i}^{2}$
$\leq\mathbb{E}[\hat{\tau}_{i}^{2}]$
where
$\hat{\tau}_{i}^{2}$ $=\left(\frac{Y_{i}(1)\mathbbm{1}\{W_{i}=1\}}{p_{i}^{MAD}(1)}-\frac{Y_{i}(0)\mathbbm{1}\{W_{i}=0\}}{p_{i}^{MAD}(0)}\right)^{2}$
$\leq\frac{Y_{i}(1)^{2}}{(\delta_{i}/2)^{2}}+\frac{Y_{i}(0)^{2}}{(\delta_{i}/2)^{2}}$
$=8M^{2}\frac{1}{\delta_{i}^{2}}$
$=8M^{2}o(i^{1/2})$
Hence,
$\sum_{i=1}^{t}\frac{\mathbb{E}[u_{i}^{2}]}{i^{2}}\leq\sum_{i=1}^{t}\frac{8M^{2}o(i^{1/2})}{i^{2}}<\infty$
by the p-series test.
Hence, by the SLLN for martingale difference sequences (Theorem 1 of Csörgő (1968)), we can conclude that
$\frac{1}{t}\sum_{i=1}^{t}u_{i}\stackrel{{\scriptstyle a.s.}}{{\to}}0.$ (8)
By Theorem 1, we have that $\hat{V}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}0$.
Therefore, applying Slutsky’s Theorem, we conclude that
$\hat{\bar{{\tau}}}_{t}+\hat{V}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}c,$ (9)
and
$\hat{\bar{{\tau}}}_{t}-\hat{V}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}c.$ (10)
So,
$\mathbb{P}\left(0\notin\left(\hat{\bar{{\tau}}}_{t}\pm\hat{V}_{t}\right)\right)$
$=\mathbb{P}\left(0<\hat{\bar{{\tau}}}_{t}-\hat{V}_{t}\right)+\mathbb{P}\left(0>\hat{\bar{{\tau}}}_{t}+\hat{V}_{t}\right)$
$\to 1,$
where the last line follows because $\mathbb{P}\left(0>\hat{\bar{{\tau}}}_{t}+\hat{V}_{t}\right)\to 0$ and $\mathbb{P}\left(0<\hat{\bar{{\tau}}}_{t}-\hat{V}_{t}\right)\to 1$.
Let $p_{t}=\mathbb{P}\left(0\in(\hat{\bar{{\tau}}}_{t}\pm\hat{V}_{t})\right)$. Since $p_{t}\to 0$ as $t\to\infty$, there exists a subsequence $t_{k}$ such that $p_{t_{k}}\leq\frac{1}{k^{2}}$, for all $k\geq 1$. Let $A_{t_{k}}$ be the event $\{0\in(\frac{1}{t_{k}}\sum_{i=1}^{t_{k}}\hat{\tau}_{i}\pm\hat{V}_{t_{k}})\}$. Then, $\sum_{k=1}^{\infty}\mathbb{P}\left(A_{t_{k}}\right)<\infty$, and by the Borel-Cantelli lemma,
$\mathbb{P}\left(\limsup_{k\to\infty}A_{t_{k}}\right)=0.$ (11)
Hence,
$\mathbb{P}\left(T_{MAD}=\infty\right)\leq\mathbb{P}\left(\limsup_{k\to\infty}A_{t_{k}}\right)=0.$ (12)
Appendix C The MAD for $K\geq 2$ Treatments
We first formalize the problem setting for $K\geq 2$ treatments, which we call the Generalized Mixture Adaptive Design, and prove more general versions of Theorems 1 and Theorem 2 for $K\geq 2$ treatments. Assume $W_{t}\in\mathcal{W}\coloneqq\{0,...,K-1\}$.
Let $\mathcal{F}_{t,n}$ be the sigma-algebra that contains all potential outcomes $\{(Y_{i}(w))_{w\in\mathcal{W}}\}_{i=1}^{t}$ and all observed data $\{W_{i},Y_{i}\}_{i=1}^{n}$ where $n\leq t$.
As before, let the assignment probabilities for any user-chosen adaptive algorithm be denoted as $p_{i}^{\mathcal{A}}(w)=\mathbb{P}(W_{t}=w\mid H_{t-1})$ where $H_{t-1}=\{W_{i},Y_{i}\}_{i=1}^{t-1}$. Hence, the Generalized Mixture Adaptive Design has assignment probabilities $p^{\text{MAD}}_{i}(w)=\delta_{i}\frac{1}{K}+(1-\delta_{i})p_{i}^{\mathcal{A}}(w)$, for all $w\in\mathcal{W}$.
For any pair of treatments $w,w^{\prime}\in\mathcal{W}$, let $\tau_{i}(w,w^{\prime})=Y_{i}(w)-Y_{i}(w^{\prime})$ and define the Average Treatment Effect (ATE) between $w$ and $w^{\prime}$ up to $t$ as $\bar{\tau}_{t}(w,w^{\prime})\coloneqq\frac{1}{t}\sum_{i=1}^{t}\tau_{i}(w,w^{\prime})$. So, using the Generalized Mixture Adaptive Design, our corresponding estimator for the ATE is:
$\hat{\bar{\tau}}_{t}(w,w^{\prime})=\frac{1}{t}\sum_{i=1}^{t}\hat{\tau_{i}}(w,w^{\prime}),$
where $\hat{\tau_{i}}(w,w^{\prime})\coloneqq\frac{\mathbbm{1}\{W_{i}=w\}Y_{i}(w)}{p^{\text{MAD}}_{i}(w)}-\frac{\mathbbm{1}\{W_{i}=w^{\prime}\}Y_{i}(w^{\prime})}{p^{MAD}_{i}(w^{\prime})}$. We also have the corresponding upper bound on the variance:
$Var(\hat{\tau}_{i}(w,w^{\prime})\mid\mathcal{F}_{i})\leq\sigma^{2}_{i}(w,w^{\prime})\text{, where }\sigma^{2}_{i}(w,w^{\prime})\coloneqq\frac{Y_{i}(w)^{2}}{p^{\text{MAD}}_{i}(w)}+\frac{Y_{i}(w^{\prime})^{2}}{p^{MAD}_{i}(w^{\prime})},$
and the analogous unbiased estimator of $\sigma^{2}_{i}(w,w^{\prime})$:
$\hat{\sigma}^{2}_{i}(w,w^{\prime})\coloneqq\frac{Y_{i}(w)^{2}\mathbbm{1}\{W_{i}=w\}}{(p^{\text{MAD}}_{i}(w))^{2}}+\frac{Y_{i}(w^{\prime})^{2}\mathbbm{1}\{W_{i}=w^{\prime}\}}{(p^{MAD}_{i}(w^{\prime}))^{2}}.$
Finally, let $S_{t}(w,w^{\prime})\coloneqq\sum_{i=1}^{t}\sigma_{i}^{2}(w,w^{\prime})$ and $\hat{S}_{t}(w,w^{\prime})\coloneqq\sum_{i=1}^{t}\hat{\sigma}_{i}^{2}(w,w^{\prime})$.
Define $T_{MAD}(w,w^{\prime})\coloneqq\inf_{t}\left\{t:0\notin(\hat{\bar{\tau}}_{t}(w,w^{\prime})\pm\hat{V}_{t}(w,w^{\prime}))\right\}$ where the treatment assignments are generated via the Generalized Mixture Adaptive Design.
Though an analogous result of Theorem 1 for this setting follows almost directly from the fact that we still have $1/\delta_{i}=o(i^{1/4})$, we provide a full statement of the result and its proof here for completeness.
For precision, we need to state our cumulative conditional variance condition for the $K\geq 2$ setting.
Assumption 3 (At Least Linear Rate of Cumulative Conditional Variances for All Pairs of Treatments).
For all $w,w^{\prime}\in\mathcal{W}$, $\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}(w,w^{\prime})\mid\mathcal{F}_{i})=\Omega(t).$
We also establish the following lemma, which we will use in the proof of Theorem 1.
Lemma A.2.
Let $w,w^{\prime}\in\{0,1,...,K-1\}$. Let $\{\hat{\tau}_{t}(w,w^{\prime})\}_{t=1}^{\infty}$ be a sequence of random variables where $W_{t}=w$ with probability $p^{\text{MAD}}_{t}(w)=\delta_{t}\frac{1}{K}+(1-\delta_{t})p_{t}^{\mathcal{A}}(w)$, for all $w\in\mathcal{W}$ and $\delta_{t}\in(0,1]$ such that $\delta_{t}=\omega\left(\frac{1}{t^{1/4}}\right)$. Assume Assumptions 1 and 3 hold. Then, $\{\hat{\tau}_{t}(w,w^{\prime})\}_{i=1}^{\infty}$ satisfies the Lindeberg-type uniform integrability condition of Waudby-Smith et al. (2023), i.e., there exists $\kappa\in(0,1)$ such that
[…]
where $B_{t}(w,w^{\prime})=\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}(w,w^{\prime})\mid\mathcal{F}_{i})$.
Proof.
By Assumption 1, for all $t,w,w^{\prime}$,
$\left|\hat{{\tau}}_{t}(w,w^{\prime})\right|=\left|\frac{Y_{t}(w)\mathbbm{1}\{W_{t}=w\}}{p^{\text{MAD}}_{t}(w)}-\frac{Y_{t}(w^{\prime})\mathbbm{1}\{W_{t}=w^{\prime}\}}{p^{\text{MAD}}_{t}(w^{\prime})}\right|\leq\frac{2M}{\min(p^{\text{MAD}}_{t}(w^{\prime}),p^{\text{MAD}}_{t}(w))}$
and $|{\tau}_{t}(w,w^{\prime})|\leq 2M$.
Hence,
$(\hat{{\tau}}_{t}(w,w^{\prime})-{\tau}_{t}(w,w^{\prime}))^{2}$
$\leq\left(\frac{2M}{\min(p^{\text{MAD}}_{t}(w),p^{\text{MAD}}_{t}(w^{\prime}))}\right)^{2}+(2M)^{2}+2\left(\frac{2M}{\min(p^{\text{MAD}}_{t}(w),p^{\text{MAD}}_{t}(w^{\prime}))}\right)2M$
$\leq\frac{24M^{2}}{(\min(p^{\text{MAD}}_{t}(w),p^{\text{MAD}}_{t}(w^{\prime})))^{2}}.$
First, note that for all $w\in\mathcal{W}$, $p^{\text{MAD}}_{t}(w)\geq\delta_{t}(1/K)$ and so $\frac{1}{p^{\text{MAD}}_{t}(w)}=o(t^{1/4})$. So, $(\hat{{\tau}}_{t}(w,w^{\prime})-{\tau}_{t}(w,w^{\prime}))^{2}=o(t^{1/2})$ almost surely.
By Assumption 2, $B_{t}(w,w^{\prime})=\Omega(t)$. Therefore, $B_{t}^{\kappa}(w,w^{\prime})=\Omega(t^{\kappa})$. Set $\kappa>1/2$. Then, there exists some $\tilde{t}$ such that for all $t\geq\tilde{t}$, $B_{t}^{\kappa}(w,w^{\prime})>(\hat{{\tau}}_{t}(w,w^{\prime})-{\tau}_{t}(w,w^{\prime}))^{2}$ a.s.. Hence, for all $t\geq\tilde{t}$, $\mathbbm{1}\{(\hat{{\tau}}_{t}(w,w^{\prime})-{\tau}_{t}(w,w^{\prime}))^{2}>(B_{t})^{\kappa}(w,w^{\prime})\}=0$, and so,
[…]
Theorem 1.
For $w,w^{\prime}\in\{0,1,...,K-1\}$, let $(\hat{\tau}_{i}(w,w^{\prime}))_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)=\frac{1}{K}\delta_{i}+(1-\delta_{i})p^{\mathcal{A}}_{i}(w)$, and $\delta_{i}\in(0,1]$ such that $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 3 hold. Then $(\hat{\bar{{\tau}}}_{t}(w,w^{\prime})\pm\hat{V}_{t}(w,w^{\prime}))$ where
$\hat{V}_{t}(w,w^{\prime})=\sqrt{\frac{2(\hat{S}_{t}(w,w^{\prime})\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}(w,w^{\prime})\eta^{2}+1}}{\alpha}\right)}$
is a valid $(1-\alpha)$ asymptotic CS for $\bar{\tau}_{t}(w,w^{\prime})$ and $\hat{V}_{t}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}0$.
Proof.
By Assumptions 1 and 3 imply that Lemma A.2 holds. Lemma A.2 and Assumption 3 satisfy Conditions L-1 and L-2 of Theorem 2.5 in Waudby-Smith et al. (2023), so, by Steps 1 and 2 of the proof of Theorem 2.5 in Waudby-Smith et al. (2023),
$(\hat{\bar{{\tau}}}_{t}(w,w^{\prime})\pm V^{*}_{t}(w,w^{\prime}))$ where
$V^{*}_{t}(w,w^{\prime})=\sqrt{\frac{2(B_{t}(w,w^{\prime})\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{B_{t}(w,w^{\prime})\eta^{2}+1}}{\alpha}\right)}$
is a valid $(1-\alpha)$asymptotic CS and $B_{t}(w,w^{\prime})=\sum_{i=1}^{t}Var(\hat{{\tau}}_{i}(w,w^{\prime})\mid\mathcal{F}_{i})$.
As noted in Equation (3),
$Var(\hat{\tau}_{i}(w,w^{\prime})\mid\mathcal{F}_{i})\leq\sigma^{2}_{i}\text{, where }\sigma^{2}_{i}\coloneqq\frac{Y_{i}(w)^{2}}{p^{\text{MAD}}_{i}(w)}+\frac{Y_{i}(w^{\prime})^{2}}{p^{MAD}_{i}(w^{\prime})}.$ (13)
Hence, $(\hat{\bar{{\tau}}}_{t}(w,w^{\prime})\pm\tilde{V}_{t}(w,w^{\prime}))$ where
$\tilde{V}_{t}(w,w^{\prime})=\sqrt{\frac{2(S_{t}(w,w^{\prime})\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{S_{t}(w,w^{\prime})\eta^{2}+1}}{\alpha}\right)}$
is still a valid $(1-\alpha)$asymptotic CS where recall, $S_{t}(w,w^{\prime})=\sum_{i=1}^{t}\sigma_{i}^{2}(w,w^{\prime})$. This holds because replacing $B_{t}(w,w^{\prime})$ with $S_{t}(w,w^{\prime})$ only makes the CS wider since $S_{t}(w,w^{\prime})\geq B_{t}(w,w^{\prime})$.
As noted in Equation (4), an unbiased estimator for $\sigma^{2}_{i}(w,w^{\prime})$ is
$\hat{\sigma}^{2}_{i}(w,w^{\prime})\coloneqq\frac{Y_{i}(w)^{2}\mathbbm{1}\{W_{i}=w\}}{(p^{\text{MAD}}_{i}(w))^{2}}+\frac{Y_{i}(w^{\prime})^{2}\mathbbm{1}\{W_{i}=w^{\prime}\}}{(p^{MAD}_{i}(w^{\prime}))^{2}}.$ (14)
and we define $\hat{S}_{t}(w,w^{\prime})=\sum_{i=1}^{t}\hat{\sigma}_{i}^{2}(w,w^{\prime})$.
To establish the validity of the CS in Theorem 1, we must show that $\frac{1}{t}\hat{S}_{t}(w,w^{\prime})-\frac{1}{t}S_{t}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}0$. Then, Condition L-3 of Theorem 2.5 of Waudby-Smith et al. (2023) is satisfied and we can conclude that $(\hat{\bar{{\tau}}}_{t}(w,w^{\prime})\pm\hat{V}_{t}(w,w^{\prime}))$ where
$\hat{V}_{t}(w,w^{\prime})=\sqrt{\frac{2(\hat{S}_{t}(w,w^{\prime})\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}(w,w^{\prime})\eta^{2}+1}}{\alpha}\right)}$
is still a valid $(1-\alpha)$ asymptotic CS.
First, note that
$(\hat{\sigma}_{i}^{2}(w,w^{\prime}))^{2}$ $\leq\left(\frac{M^{2}}{p_{i}^{MAD}(w)^{2}}+\frac{M^{2}}{p_{i}^{MAD}(w^{\prime})^{2}}\right)^{2}$
$=\frac{M^{4}}{p_{i}^{MAD}(w)^{4}}+\frac{M^{4}}{p_{i}^{MAD}(w^{\prime})^{4}}+2\frac{M^{2}}{p_{i}^{MAD}(w)^{2}p_{i}^{MAD}(w^{\prime})^{2}}$
$\leq\frac{M^{4}}{(\delta_{i}/K)^{4}}+\frac{M^{4}}{(\delta_{i}/K)^{4}}+2\frac{M^{2}}{(\delta_{i}/K)^{4}}$
$=M^{4}\left(K^{4}\frac{1}{\delta_{i}^{4}}+K^{4}\frac{1}{\delta_{i}^{4}}+K^{5}\frac{1}{\delta_{i}^{4}}\right)$
$=o(i)$
where the last line follows because $\frac{1}{\delta_{i}}=o\left(i^{1/4}\right)$ as shown in Lemma A.1. Define $X_{i}(w,w^{\prime})=\hat{\sigma}_{i}^{2}(w,w^{\prime})-\sigma_{i}^{2}(w,w^{\prime})$ and note that $X_{i}(w,w^{\prime})$ is a martingale difference sequence. Hence,
$\mathbb{E}[X_{i}(w,w^{\prime})]$ $=E[(\hat{\sigma}_{i}(w,w^{\prime})^{2}]-(\sigma_{i}(w,w^{\prime})^{2}$
$\leq E[(\hat{\sigma}_{i}(w,w^{\prime})^{2}]$
$=o(i)$
So, $\frac{\mathbb{E}[X_{i}^{2}(w,w^{\prime})]}{i^{2}}=\frac{o(i)}{i^{2}}$ and hence, $\sum_{i=1}^{\infty}\frac{\mathbb{E}[X_{i}^{2}(w,w^{\prime})]}{i^{2}}<\infty$. For instance, if $\delta_{i}=\frac{1}{i^{a}}$ for some $40\leq a<1/$ 4, then $\frac{\mathbb{E}[X_{i}^{2}(w,w^{\prime})]}{i^{2}}\leq Ci^{4a-2}$ for some constant $C<\infty$, and since $14a-2<-$ 1, $\sum_{i=1}^{\infty}\frac{\mathbb{E}[X_{i}^{2}(w,w^{\prime})]}{i^{2}}\leq C\sum_{i=1}^{\infty}i^{4a-2}<\infty$ by the p-series test.
Then, by the SLLN for martingale difference sequences (Theorem 1 of Csörgő (1968)), we can conclude that
$\frac{1}{t}\sum_{i=1}^{t}X_{i}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}0.$
Hence, Condition L-3 of Waudby-Smith et al. (2023) is satisfied and by Step 3 of the proof for Theorem 2.5 of Waudby-Smith et al. (2023), we can conclude that $(\hat{\bar{{\tau}}}_{t}(w,w^{\prime})\pm\hat{V}_{t}(w,w^{\prime}))$ where
$\hat{V}_{t}(w,w^{\prime})=\sqrt{\frac{2(\hat{S}_{t}(w,w^{\prime})\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}(w,w^{\prime})\eta^{2}+1}}{\alpha}\right)}$
is still a valid $(1-\alpha)$ asymptotic CS.
To show $\hat{V}_{t}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}0$, note that
$\hat{\sigma}_{i}^{2}$ $\leq\frac{M^{2}}{p_{i}^{MAD}(w)^{2}}+\frac{M^{2}}{p_{i}^{MAD}(w^{\prime})^{2}}$
$\leq\frac{M^{2}}{(\delta_{i}/K)^{2}}+\frac{M^{2}}{(\delta_{i}/K)^{2}}$
$=\frac{K^{2}M^{2}}{\delta_{i}^{2}}+\frac{K^{2}M^{2}}{\delta_{i}^{2}}$
$=2K^{2}M^{2}\frac{1}{\delta_{i}^{2}}$
$=2K^{2}M^{2}o(i^{1/2})$
So, $\hat{S}_{t}(w,w^{\prime})\leq 8M^{2}\sum_{i=1}^{t}o(i^{1/2})$ a.s..
So, there exists positive real numbers $N$ and $x_{0}$ such that for all $t\geq x_{0}$, $\hat{S}_{t}(w,w^{\prime})\leq N\sum_{i=1}^{t}i^{1/2}<N\sum_{i=1}^{t}t^{1/2}=Nt^{1+1/2}$, and hence, we can conclude that $\hat{S}_{t}(w,w^{\prime})=O(t^{1+\frac{1}{2}})$.
We can show that $\log\left(\hat{S}_{t}(w,w^{\prime})\right)=o(t^{1/2})$ a.s.. For $t\geq x_{0}$,
$\frac{\log(\hat{S}_{t}(w,w^{\prime}))}{t^{1/2}}\leq\frac{\log(Nt^{3/2})}{t^{1/2}}=\frac{\log(N)+(3/2)\log(t))}{t^{1/2}}$
and $\frac{\log(N)+(3/2)\log(t))}{t^{1/2}}\to 0$ as $t\to\infty$. Therefore, $\log(\hat{S}_{t}(w,w^{\prime}))=o(t^{1/2})$ a.s. and $\hat{S}_{t}(w,w^{\prime})\log(\hat{S}_{t}(w,w^{\prime}))=o(t^{2})$ a.s.. Hence, $\hat{V}_{t}(w,w^{\prime})=\frac{2(\hat{S}_{t}(w,w^{\prime})\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}_{t}(w,w^{\prime})\eta^{2}+1}}{\alpha}\right)=o(1)$ a.s..
Note, if $1/\delta_{i}$ was increasing at a rate faster than $i^{1/4}$ asymptotically, e.g., $\delta_{i}=O(i^{1/2})$, we are not guaranteed that $\hat{V}_{t}(w,w^{\prime})=o(1)$ a.s.. ∎
Theorem 2.
For $w,w^{\prime}\in\{0,1,...,K-1\}$, let $(\hat{\tau}_{i}(w,w^{\prime}))_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)=\delta_{i}\left(\frac{1}{2}\right)+(1-\delta_{i})p^{\mathcal{A}}_{i}(w)$ for $w\in\mathcal{W}$ and $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 3 hold. Assume $\bar{\tau_{i}}(w,w^{\prime})\to c$ for some $|c|>0$ as $i\to\infty$. Then,
$\mathbb{P}\left(T_{MAD}(w,w^{\prime})<\infty\right)=1.$
Proof.
We will first show that $\frac{1}{t}\sum_{i=1}^{t}\hat{\tau}_{i}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}c$. Let $u_{i}(w,w^{\prime})\coloneqq\hat{\tau}_{i}(w,w^{\prime})-\tau_{i}(w,w^{\prime})$.
Note that
$\mathbb{E}[u_{i}(w,w^{\prime})^{2}]$ $=\mathbb{E}[\hat{\tau}_{i}(w,w^{\prime})^{2}]-\tau_{i}(w,w^{\prime})^{2}$
$\leq\mathbb{E}[\hat{\tau}_{i}(w,w^{\prime})^{2}]$
where
$\hat{\tau}_{i}(w,w^{\prime})^{2}$ $=\left(\frac{Y_{i}(w)\mathbbm{1}\{W_{i}=w\}}{p_{i}^{MAD}(w)}-\frac{Y_{i}(w^{\prime})\mathbbm{1}\{W_{i}=w^{\prime}\}}{p_{i}^{MAD}(w^{\prime})}\right)^{2}$
$\leq\frac{Y_{i}(w)^{2}}{(\delta_{i}/K)^{2}}+\frac{Y_{i}(w^{\prime})^{2}}{(\delta_{i}/K)^{2}}$
$=2K^{2}M^{2}\frac{1}{\delta_{i}^{2}}$
$=2K^{2}M^{2}o(i^{1/2})$
Hence,
$\sum_{i=1}^{t}\frac{\mathbb{E}[u_{i}(w,w^{\prime})^{2}]}{i^{2}}\leq\sum_{i=1}^{t}\frac{8M^{2}o(i^{1/2})}{i^{2}}<\infty$
by the p-series test. Hence, by the SLLN for martingale difference sequences (Theorem 1 of Csörgő (1968)), we can conclude that
$\frac{1}{t}\sum_{i=1}^{t}u_{i}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}0.$
By Theorem 1, we have that $\hat{V}_{t}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}0$.
Therefore, applying Slutsky’s Theorem, we conclude that
$\hat{\bar{{\tau}}}_{t}(w,w^{\prime})+\hat{V}_{t}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}c,$
and
$\hat{\bar{{\tau}}}_{t}(w,w^{\prime})-\hat{V}_{t}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}c.$
So,
$\mathbb{P}\left(0\notin\left(\hat{\bar{{\tau}}}_{t}(w,w^{\prime})\pm\hat{V}_{t}(w,w^{\prime})\right)\right)$
$=\mathbb{P}\left(0<\hat{\bar{{\tau}}}_{t}(w,w^{\prime})-\hat{V}_{t}(w,w^{\prime})\right)+\mathbb{P}\left(0>\hat{\bar{{\tau}}}_{t}(w,w^{\prime})+\hat{V}_{t}(w,w^{\prime})\right)$
$\to 1,$
where the last line follows because $\mathbb{P}\left(0>\hat{\bar{{\tau}}}_{t}(w,w^{\prime})+\hat{V}_{t}(w,w^{\prime})\right)\to 0$ and
$\mathbb{P}\left(0<\hat{\bar{{\tau}}}_{t}(w,w^{\prime})-\hat{V}_{t}(w,w^{\prime})\right)\to 1$.
Let $p_{t}=\mathbb{P}\left(0\in(\hat{\bar{{\tau}}}_{t}(w,w^{\prime})\pm\hat{V}_{t}(w,w^{\prime}))\right)$. Since $p_{t}\to 0$ as $t\to\infty$, there exists a subsequence $t_{k}$ such that $p_{t_{k}}\leq\frac{1}{k^{2}}$, for all $k\geq 1$. Let $A_{t_{k}}$ be the event $\{0\in(\frac{1}{t_{k}}\sum_{i=1}^{t_{k}}\hat{\tau}_{i}(w,w^{\prime})\pm\hat{V}_{t_{k}}(w,w^{\prime}))\}$. Then, $\sum_{k=1}^{\infty}\mathbb{P}\left(A_{t_{k}}\right)<\infty$, and by the Borel-Cantelli lemma,
$\mathbb{P}\left(\limsup_{k\to\infty}A_{t_{k}}\right)=0.$
Hence,
$\mathbb{P}\left(\tilde{T}_{MAD}(w,w^{\prime})=\infty\right)\leq\mathbb{P}\left(\limsup_{k\to\infty}A_{t_{k}}\right)=0.$
Appendix D The MAD for Batched Assignment Algorithms
We now define the causal estimators and estimands for the batched setting. Let $\{(Y_{i}^{(j)}(w))_{w\in\mathcal{W}}\}$ be the set of all potential outcome for unit $i$ in batch $j$. Given a pair of treatments $w,w^{\prime}\in\mathcal{W}\coloneqq\{0,1,...,K-1\}$ for $K\geq 2$, we can define the Average Treatment Effect within a batch $j$ as:
$\tau^{\text{batched}}_{j}(w,w^{\prime})=\frac{1}{B}\sum_{i=1}^{B}Y_{i}^{(j)}(w)-Y_{i}^{(j)}(w^{\prime}),$
and the corresponding unbiased estimator as:
$\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime})=\frac{1}{B}\sum_{i=1}^{B}\hat{\tau}^{(j)}_{i}(w,w^{\prime}).$
where
$\hat{\tau}^{(j)}_{i}(w,w^{\prime})=\frac{Y_{i}^{(j)}(w)\mathbbm{1}\{W^{(j)}_{i}=w\}}{p^{\text{MAD}_{\text{batched}}}_{j}(w)}-\frac{Y^{(j)}_{i}(w^{\prime})\mathbbm{1}\{W^{(j)}_{i}=w^{\prime}\}}{p^{\text{MAD}_{\text{batched}}}_{j}(w^{\prime})}.$
Thus, our target estimand, the Batched-Average Treatment Effect up to batch $b$, is:
$\bar{\tau}^{\text{batched}}_{b}(w,w^{\prime})=\frac{1}{b}\sum_{j=1}^{b}\tau^{\text{batched}}_{j}(w,w^{\prime}).$
and the corresponding unbiased estimator is:
$\hat{\bar{\tau}}^{\text{batched}}_{b}(w,w^{\prime})=\frac{1}{b}\sum_{j=1}^{b}\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime}).$
Hence, we define $\mathcal{F}^{\text{batched}}_{b}$ as the filtration generated by all potential outcomes for all units up to and including the $b$th batch $\{\{(Y_{i}^{(j)}(w))_{w\in\mathcal{W}}\}_{i=1}^{B_{h}}\}_{j=1}^{b}$ and all observed history up to batch $b-1$, $\{H^{\text{batched}}_{j}\}_{j=1}^{b-1}$.
Then, for each $i=1,...,B$, $Var\left(\hat{\tau}^{(j)}_{i}(w,w^{\prime})\mid\mathcal{F}^{\text{batched}}_{j}\right)\leq\sigma_{i}^{(j)2}(w,w^{\prime})$ where
$\sigma_{i}^{(j)2}(w,w^{\prime})=\frac{Y_{i}(w)^{2}}{p^{\text{MAD}_{\text{batched}}}_{j}(w)}+\frac{Y_{i}(w^{\prime})^{2}}{p^{\text{MAD}_{\text{batched}}}_{j}(w^{\prime})}.$
So, $\hat{\sigma}_{i}^{(j)2}(w,w^{\prime})=\frac{Y_{i}(w)^{2}\mathbbm{1}\{W_{i}=w\}}{(p^{\text{MAD}_{\text{batched}}}_{j}(w))^{2}}+\frac{Y_{i}(w^{\prime})^{2}\mathbbm{1}\{W_{i}=w^{\prime}\}}{(p^{\text{MAD}_{\text{batched}}}_{j}(w^{\prime}))^{2}}.$
Hence,
$Var\left(\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime})\mid\mathcal{F}^{\text{batched}}_{j}\right)=\frac{1}{B^{2}}\sum_{i=1}^{B}Var\left(\hat{\tau}^{(j)}_{i}(w,w^{\prime})\mid\mathcal{F}^{\text{batched}}_{j}\right)\leq\frac{1}{B^{2}}\sum_{i=1}^{B}\sigma_{i}^{(j)2}(w,w^{\prime}).$
Hence, we define
$S^{\text{batched}}_{b}(w,w^{\prime})\coloneqq\sum_{j=1}^{b}\frac{1}{B^{2}}\sum_{i=1}^{B}\sigma_{i}^{(j)2}(w,w^{\prime})$ and $\hat{S}^{\text{batched}}_{b}(w,w^{\prime})\coloneqq\sum_{j=1}^{b}\frac{1}{B^{2}}\sum_{i=1}^{H_{j}}\hat{\sigma}_{i}^{(j)2}(w,w^{\prime})$.
Finally, we state the analogous assumptions as Assumptions 1 and 2 for the batched assignment setting.
Assumption 4 (Bounded Potential Outcomes for Batched Bandits).
There exists $M\in\mathbb{R}$ such that
$|Y^{(j)}_{i}(w)|\leq M<\infty$
for all $j,i\in\mathbb{N}^{+},w\in\mathcal{W}$.
Assumption 5 (At Least Linear Rate of Cumulative Conditional Variances for Batched Bandits).
For all $w,w^{\prime}\in\mathcal{W}$,
$\sum_{j=1}^{b}Var\left(\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime})\mid\mathcal{F}^{\text{batched}}_{j}\right)=\Omega(b).$
Theorem $1^{\dagger}$.
For $w,w^{\prime}\in\mathcal{W}$, let $(\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime}))_{j=1}^{\infty}$ be the sequence of random variables where $W^{(j)}_{i}=w$ with probability $p_{j}^{\text{MAD}_{\text{batched}}}(w)$ as in Definition 5, $w\in\mathcal{W}$, and $\delta_{j}\in(0,1]$ such that $\delta_{j}=\omega\left(\frac{1}{j^{1/4}}\right)$. Assume Assumptions 4 and 5 hold. Then $(\hat{\bar{{\tau}}}^{\text{batched}}_{b}(w,w^{\prime})\pm\hat{V}^{\text{batched}}_{b}(w,w^{\prime}))$ where
$\hat{V}^{\text{batched}}_{b}(w,w^{\prime})\coloneqq\sqrt{\frac{2(\hat{S}^{\text{batched}}_{b}\eta^{2}+1)}{t^{2}\eta^{2}}\log\left(\frac{\sqrt{\hat{S}^{\text{batched}}_{b}\eta^{2}+1}}{\alpha}\right)}$
is a valid $(1-\alpha)$ asymptotic CS for $\bar{\tau}^{\text{batched}}_{b}$ and $\hat{V}^{\text{batched}}_{b}(w,w^{\prime})\stackrel{{\scriptstyle a.s.}}{{\to}}0$.
Proof.
Note, the above result is equivalent to Theorem 1 except we replace $\hat{\tau}_{i}(w,w^{\prime})$ with $\hat{\tau}_{j}^{\text{batched}}(w,w^{\prime})$, $B_{t}(w,w^{\prime})$ with $B^{\text{batched}}_{b}(w,w^{\prime})=\sum_{j=1}^{b}Var\left(\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime})\mid\mathcal{F}^{\text{batched}}_{j,j-1}\right)$, and $\hat{S}_{t}(w,w^{\prime})$ and $S_{t}(w,w^{\prime})$ with $S^{\text{batched}}_{b}(w,w^{\prime})\coloneqq\sum_{j=1}^{b}\frac{1}{B^{2}}\sum_{i=1}^{B}\sigma_{i}^{(j)2(w,w^{\prime})}$ and $\hat{S}^{\text{batched}}_{b}(w,w^{\prime})\coloneqq\sum_{j=1}^{b}\frac{1}{B^{2}}\sum_{i=1}^{H_{j}}\hat{\sigma}_{i}^{(j)2(w,w^{\prime})}$ respectively. Assumption 5 ensures $\sum_{j=1}^{b}Var\left(\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime})\mid\mathcal{F}^{\text{batched}}_{j,j-1}\right)=\Omega(b)$ and, because $p_{j}^{\text{MAD}_{\text{batched}}}(w)\geq\frac{1}{K}\delta_{j}$, we have that $(\hat{\tau}_{j}^{\text{batched}}(w,w^{\prime})-{\tau}_{j}^{\text{batched}}(w,w^{\prime}))^{2}=o(j^{1/2})$ and following the steps of the proof of Lemma A.2, we establish that $\{\hat{\tau}_{j}^{\text{batched}}(w,w^{\prime})\}_{j=1}^{\infty}$ satisfies the Lindeberg-type uniform integrability condition of Waudby-Smith et al. (2023), i.e., there exists $\kappa\in(0,1)$ such that
[…]
Hence, the remainder of the proof follows directly from Theorem 1, replacing the corresponding terms for the batched bandit setting, because the fact that $p_{j}^{\text{MAD}_{\text{batched}}}(w)\geq\frac{1}{K}\delta_{j}$ ensures that all analogous terms have the same rates as in Theorem 1. ∎
For $w,w^{\prime}\in\mathcal{W}$, define $T_{\text{MAD}_{\text{batched}}}(w,w^{\prime})\coloneqq\inf_{t}\left\{b:0\notin(\hat{\bar{\tau}}^{\text{batched}}_{b}(w,w^{\prime})\pm V^{\text{batched}}_{b}(w,w^{\prime}))\right\}$. We now state and prove an analogous result to Theorem 2 in the batched assignment setting.
Theorem $2^{\dagger}$.
For $w,w^{\prime}\in\mathcal{W}$, let $(\hat{\tau}^{\text{batched}}_{j}(w,w^{\prime}))_{j=1}^{\infty}$ be the sequence of random variables where $W^{(j)}_{i}=w$ with probability $p_{j}^{\text{MAD}_{\text{batched}}}(w)$ as in Defintion 5, and $\delta_{j}\in(0,1]$ such that $\delta_{j}=\omega\left(\frac{1}{j^{1/4}}\right)$. Assume Assumptions 4 and 5 hold. Then, if $\hat{\bar{\tau}}^{\text{batched}}_{b}(w,w^{\prime})\to c$ as $b\to\infty$ for some $|c|>0$,
$\mathbb{P}\left(T_{\text{MAD}_{\text{batched}}}(w,w^{\prime})<\infty\right)=1.$
Proof.
Having established Theorem $1^{\dagger}$, the remainder of the proof follows from Theorem 2, replacing the corresponding terms for the batched bandit setting, because the fact that $p_{j}^{\text{MAD}_{\text{batched}}}(w)\geq\frac{1}{K}\delta_{j}$ ensures that all analogous terms have the same rates as in Theorem 2. ∎
Appendix E Proof of Theorem 3
Recall, the definition of asymptotic confidence sequence for $\hat{\bar{\tau}}_{t}$ states that $(\hat{\bar{\tau}}_{t}\pm\hat{V}_{t})$ is a two-sided $(1-\alpha)$ asymptotic CS for $\hat{\bar{\tau}}_{t}$ if there exists an (unknown) non-asymptotic confidence sequence $(\hat{\bar{\tau}}_{t}\pm V^{*}_{t})$ for $\tau_{t}$ such that $\frac{V^{*}_{t}}{\hat{V}_{t}}\stackrel{{\scriptstyle a.s.}}{{\to}}1$. Furthermore, we say that $\hat{V}_{t}$ has approximation rate $R_{t}$ if $V^{*}_{t}=\hat{V}_{t}+o(R_{t})$ almost surely, where $R_{t}$ is a deterministic (i.e., non-random) sequence.
Define $T^{*}_{MAD}\coloneqq\inf_{t}\left\{t:0\notin(\hat{\bar{\tau}}_{t}\pm V^{*}_{t})\right\}$, i.e., this is the analogous stopping time to $T_{MAD}$ for the non-asymptotic anytime valid CS.
We will show that, under the same setting as Theorem 2, $T^{*}_{MAD}<\infty$ with probability $1$.
Theorem 3.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)=\frac{1}{2}\delta_{i}+(1-\delta_{i})p^{\mathcal{A}}_{i}(w)$, $w\in\{0,1\}$, and $\delta_{i}=\omega\left(\frac{1}{i^{1/4}}\right)$. Assume Assumptions 1 and 2 hold. Assume $\bar{\tau}_{t}\to c$ for some $|c|>0$ as $i\to\infty$. Then, $\mathbb{P}\left(T^{*}_{MAD}<\infty\right)=1$.
At a high level, the proof proceeds by showing that $\delta_{t}$ guarantees that $V^{*}_{t}=\hat{V}_{t}+o(1)$ and follows almost immediately from results already established in the proofs of Theorem 1 and 2. We prove this result for the binary treatment setting, though it can be naturally extended to more than two treatments, as in Theorem 2, and the batched bandits setting, as in Theorem $2^{\dagger}$.
Proof.
As established Theorem 2.5 of Waudby-Smith et al. (2023), the asymptotic CS of Theorem 1 has approximation rate $R_{t}=\sqrt{S_{t}\log(S_{t})}/t$. For $\delta_{t}=\omega(1/t^{1/4})$, we have that
$\sqrt{S_{t}\log(S_{t})}/t\leq\sqrt{\hat{S}_{t}\log(\hat{S}_{t})}/t=o(1)$
where the final equality follows from Equation (7) in the proof of Theorem 1. Hence, for the MAD with $\delta_{i}=\omega(1/i^{1/4})$ for all $i=1,...,t$, $V^{*}_{t}=\hat{V}_{t}+o(1)$ almost surely.
By Equations (9) and (10) in the proof of Theorem 2,
$\hat{\bar{\tau}}_{t}+V^{*}_{t}$ =$\stackrel{{\scriptstyle a.s.}}{{=}}\hat{\bar{\tau}}_{t}+\hat{V}_{t}+o(1)\stackrel{{\scriptstyle a.s.}}{{\to}}c$ (15)
$\hat{\bar{\tau}}_{t}-V^{*}_{t}$ =$\stackrel{{\scriptstyle a.s.}}{{=}}\hat{\bar{\tau}}_{t}-\hat{V}_{t}-o(1)\stackrel{{\scriptstyle a.s.}}{{\to}}c.$ (16)
So, we can conclude that $\mathbb{P}\left(0\in(\hat{\bar{\tau}}_{t}-V^{*}_{t})\right)\to 1$ and the fact that $\mathbb{P}(T^{*}_{MAD}<\infty)=1$ follows immediately from Equations (11) and (12) of the proof of Theorem 2.
Appendix F Proof of Lemma 3.1
Recall, we define $\tilde{T}_{MAD}\coloneqq\inf_{t}\left\{t:0\notin(\hat{\bar{\tau}}_{t}\pm u(W_{t})/t\right\}$ where $u(W_{t})$ is defined as on Corollary 2 of Howard et al. (2021) and $W_{t}=\sum_{i=1}^{t}(\hat{\tau}_{i}-\hat{\bar{\tau}}_{i})^{2}$.
Theorem 3.
Let $(\hat{\tau}_{i})_{i=1}^{\infty}$ be the sequence of random variables where $W_{i}=w$ with probability $p_{i}^{\text{MAD}}(w)=\frac{1}{2}\delta_{i}+(1-\delta_{i})p^{\mathcal{A}}_{i}(w)$, $w\in\{0,1\}$, where there exists some $10<p_{min}<$ 1 such that $\delta_{i}\geq\min\{p_{min},1-p_{min}\}$ for all $i\in[1,\infty)$. Assume Assumptions 1 and 2 hold. Assume $\bar{\tau}_{t}\to c$ for some $|c|>0$ as $i\to\infty$. Then, $\mathbb{P}\left(\tilde{T}_{MAD}<\infty\right)=1$.
Proof.
Because $\delta_{i}\geq\min\{p_{min},1-p_{min}\}$ for all $i\in[1,\infty)$, $p_{i}^{\text{MAD}}(w)\geq\frac{p_{min}}{2}$. By Theorem 1 of Howard et al. (2021), $u(W_{t})/t$ is guaranteed to shrink to zero as $t$ grows. By Equation (8) of the proof of Theorem 2, we have that $\hat{\bar{\tau}}_{t}\stackrel{{\scriptstyle a.s.}}{{\to}}c$.
Therefore, applying Slutsky’s Theorem, we conclude that
$\hat{\bar{{\tau}}}_{t}+u(W_{t})/t\stackrel{{\scriptstyle a.s.}}{{\to}}c,$ (17)
and
$\hat{\bar{{\tau}}}_{t}-u(W_{t})/t\stackrel{{\scriptstyle a.s.}}{{\to}}c.$ (18)
and the stopping time result follows immediately from the Borel Cantelli argument of Equations (11) and (12) of the proof of Theorem 2.
Hence, choosing $\delta_{i}$ as we did for the results of Section 7 enables the experimenter to have a guaranteed stopping time result using the MAD with a non-asymptotic CS. However, we find empirically that MAD with the asymptotic CS of Theorem 1 often achieves the correct coverage in finite samples while shrinking significantly faster than the non-asymptotic CS of Howard et al. (2021), and hence, we recommend the use of the asymptotic CS of Theorem 1 in practice.
Appendix G Additional Simulation Results
We repeat versions of the Bernoulli outcome experiments in Section 7 with different outcome models. As in the Bernoulli experiments, we consider both a Thompson sampler and UCB and run all experiments across $N=100$ random seeds. As baselines, we use the Standard Bandit design and the Bernoulli design. MAD is implemented with $\delta(t)=\frac{1}{t^{0.24}}$. First, we consider a two-arm bandit with a Normal outcome model, i.e., $Y_{i}(1)\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(\mu_{0},1)$ and $Y_{i}(0)\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathcal{N}(\mu_{1},1)$, so $\bar{\tau}_{t}=\mu_{1}-\mu_{0}$ for all $t$. The Thompson sampler is implemented with independent standard Normal priors on both arms. We consider $(\mu_{0},\mu_{1})=\{(1,1),(1,2),(1,4)\}$ so the ATE = $0,1,3$, respectively. We set $T=1000$.
Figures 5– 8 show the average width, coverage, stopping behavior, and time averaged reward of this experiment, respectively. Recall Section 7 for descriptions of how each of these metrics are computed. As in the Bernoulli experiments, the Standard Bandit design does not have anytime validity. Though it can have smaller width than the MAD and the Bernoulli designs, the CSs generated via the Standard Bandit design cannot be used to produce valid inference. On the other hand, the MAD achieves the proper anytime validity in finite-samples and performs similarly to the Bernoulli design in terms of both CS width and stopping time while having reward closer to that of the Standard Bandit design.
Figure 5: Average value of $\hat{V}_{t}$, which defines the width of the CS of Theorem 1, across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Normal outcome model, as described in Appendix G. Error bands depict ±$\pm 2$ SEs.
Figure 6: Empirical coverage of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Normal outcome model, as described in Appendix G. $1-\alpha$ is depicted by the dotted black line. Error bands depict ±$\pm 2$ SEs.
Figure 7: The proportion of replicates which do not contain 00 at time $t$ of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Normal outcome model, as described in Appendix G. Error bands depict ±$\pm 2$ SEs.
Figure 8: The time averaged reward of each of the designs across $N=100$ random seeds under a two-armed bandit setting with a Normal outcome model, as described in Section G. Error bands depict ±$\pm 2$ SEs.
We also consider misspecified settings where the Thompson sampler still assumes a Normal outcome model with standard normal priors as in the previous experiment, but the outcome model is not normally distributed. Instead, we set the outcome model to be t-distributed or Cauchy, i.e., $Y_{i}(1)\stackrel{{\scriptstyle i.i.d.}}{{\sim}}t(\mu_{0},\nu)$ and $Y_{i}(0)\stackrel{{\scriptstyle i.i.d.}}{{\sim}}t(\mu_{1},\nu)$, where $\nu$ is the degrees-of-freedom ($\nu=1$ corresponds to the Cauchy distribution), and $\mu_{i}$ is the non-centrality parameter, $i=0,1$. In these misspecified settings, MAD is implemented with $\delta(t)=\frac{1}{t^{0.2}}$ to better control the variance of the ATE estimator. As in the Normal example, we set $T=1000$ and run all experiments across $N=100$ random seeds. Figures 9– 12 shows the corresponding results across different choices of $\nu$ for t-distributed outcomes with $(\mu_{0},\mu_{1})=(1,2)$. Figures 9– 12 shows the corresponding results for the Cauchy outcome distribution across $(\mu_{0},\mu_{1})=\{(1,1),(1,2),(1,4)\}$. Note, this is the same set of $(\mu_{0},\mu_{1})$ pairs as used in the Normal outcome distribution examples.
Notably, even under heavy-tailed t-distributed outcomes (e.g., $\nu=5$), the MAD still performs similarly to the Bernoulli design in terns of both width, coverage, and stopping times, while still having reward closer to the Standard Bandit design (which once again, does not maintain anytime validity in this setting). However, under a Cauchy outcome model, as shown in Figures 13– 16, the width and reward of all designs, even the Bernoulli, can be unstable due to the heavy tails of the Cauchy distribution, and when the ATE is large, the MAD can sometimes not have the correct coverage. However, we note that the ATE is not well-defined in this setting since the expectation of a Cauchy is undefined, and hence, the IPW estimator is no longer an unbiased estimator for the true ATE and no method is guaranteed to maintain validity in this setting.
Finally, we assess all designs using the asymptotic CS of Theorem 1 and the non-asymptotic CS of Howard et al. (2021) in the Bernoulli outcome setting described in Section 7 to exhibit how the MAD with our asymptotic CS generally produces far more powerful inference than the MAD used with the non-asymptotic CS of Howard et al. (2021) while still maintaining finite-sample anytime validity; hence, we recommend the use of our asymptotic CS in practice. In this setting, we use the MAD setting $\delta_{t}$ like in Example 2 of Section 6, as the CS of Howard et al. (2021) requires there exists some $p_{min}$ such that all treatment assignment probabilities are bounded within $[p_{min},1-p_{min}]$. Because of this restriction, we can only use Thompson sampling with the CS of Howard et al. (2021). We also implement the Standard Bandit and Bernoulli designs and assess the Howard et al. (2021) CS for them. Since $p_{min}$ does not exist for the Standard Bandit design with Thompson sampling, we set $p_{min}$ heuristically to be $1/T$.  [^2] Since we expect Thompson sampling to achieve a regret of $O(\log(T))$ asymptotically Agrawal and Goyal (2012), the share of units that get assigned to the suboptimal treatment is on the order of $\log(T)/T$, so we use $1/T$ as a conservative, heuristic lower bound. Note, however that the CS of Howard et al. (2021) is not guaranteed to have error control for the Standard Bandit design because their validity results require the existence of such a $p_{min}$. As shown in Figures 17– 19, the CS of Howard et al. (2021) is generally much wider than that of our asymptotic CS and the stopping times occur much later for the MAD. Since our asymptotic CS still maintains the proper error control in finite samples, we recommend its use in practice.
Figure 9: Average value of $\hat{V}_{t}$, which defines the width of the CS of Theorem 1, across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a t-distributed outcome model, as described in Section G. Note, $(\mu_{0},\mu_{1})=(1,2)$ and hence the ATE$=1$ for all experiments depicted here. Error bands depict ±$\pm 2$ SEs.
Figure 10: Empirical coverage of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a t-distributed outcome model, as described in Section G. $1-\alpha$ is depicted by the dotted black line. Note, $(\mu_{0},\mu_{1})=(1,2)$ and hence the ATE$=1$ for all experiments depicted here. Error bands depict ±$\pm 2$ SEs.
Figure 11: The proportion of replicates which do not contain 00 at time $t$ of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a t-distributed outcome model, as described in Appendix G. Note, $(\mu_{0},\mu_{1})=(1,2)$ and hence the ATE$=1$ for all experiments depicted here. Error bands depict ±$\pm 2$ SEs.
Figure 12: The time averaged reward of each of the designs across $N=100$ random seeds under a two-armed bandit setting with a t-distributed outcome model, as described in Appendix G. Note, $(\mu_{0},\mu_{1})=(1,2)$ and hence the ATE$=1$ for all experiments depicted here. Error bands depict ±$\pm 2$ SEs.
Figure 13: Average value of $\hat{V}_{t}$, which defines the width of the CS of Theorem 1, across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Cauchy outcome model, as described in Appendix G. Error bands depict ±$\pm 2$ SEs.
Figure 14: Empirical coverage of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Cauchy outcome model, as described in Appendix G. $1-\alpha$ is depicted by the dotted black line. Error bands depict ±$\pm 2$ SEs.
Figure 15: The proportion of replicates which do not contain 00 at time $t$ of the CS of Theorem 1 across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Cauchy outcome model, as described in Appendix G. $1-\alpha$ is depicted by the dotted black line. Error bands depict ±$\pm 2$ SEs.
Figure 16: The time averaged reward of each of the designs across $N=100$ random seeds under a two-armed bandit setting with a Cauchy outcome model, as described in Appendix G. Error bands depict ±$\pm 2$ SEs.
Figure 17: Average width of the CS of Theorem 1 vs. that of Howard et al. (2021) across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Bernoulli outcome model, as described in Section 7. Error bands depict ±$\pm 2$ SEs.
Figure 18: Empirical coverage of the CS of Theorem 1 vs. that of Howard et al. (2021) across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Bernoulli outcome model, as described in Section 7. $1-\alpha$ is depicted by the dotted black line. Error bands depict ±$\pm 2$ SEs.
Figure 19: The proportion of replicates which do not contain 00 at time $t$ of the CS of Theorem 1 vs. that of Howard et al. (2021) across $N=100$ random seeds for different experimental designs under a two-armed bandit setting with a Bernoulli outcome model, as described in Section 7. Error bands depict ±$\pm 2$ SEs.
[…]